### v1

# ChatCFD: an End-to-End CFD Agent with Domain-specific Structured Thinking

Abstract
Computational Fluid Dynamics (CFD) is critical for scientific and engineering advancements but is hindered by operational complexity and the need for extensive domain expertise. This paper presents ChatCFD, a large language model (LLM)-driven pipeline that automates computational fluid dynamics (CFD) workflows within the OpenFOAM framework, enabling users to configure and execute complex simulations from natural language prompts or published literature with minimal prior expertise. A key innovation is its ability to think structurally in database construction, configuration validation, and error reflection, which systematically integrates domain-specific CFD and OpenFOAM knowledge (e.g., solvers, turbulence models, file dependencies) with general LLMs, thereby enhancing accuracy and adaptability. More specifically, ChatCFD employs a four-stage workflow: (1) Knowledge Base Construction, creating a structured JSON database from OpenFOAM tutorials and manuals; (2) User Input Processing, guiding users via a multimodal interface (dialogue, documents, mesh files); (3) Case File Initialization, generating case files using a preprocessed knowledge base; and (4) Simulation Execution and Error Correction, running simulations and resolving errors with Retrieval-Augmented Generation (RAG). Powered by DeepSeek-R1 and DeepSeek-V3 models, a multi-agent architecture, and specialized OpenFOAM knowledge, ChatCFD features an interactive chat interface, hierarchical parameter extraction, and a robust error correction system addressing dimension mismatches, missing files, persistent errors, and general issues. Validation results demonstrate that ChatCFD autonomously reproduces published CFD literature results without human intervention, a challenging task involving complex, unseen configurations beyond basic CFD examples and general LLM capabilities. This achievement underscores the significant advancement enabled by its domain-specific structured thinking ability. In addition, validation experiments demonstrate error-free configurations in 30-40% of incompressible and compressible CFD cases and operational success in 60-80%, establishing a benchmark for automated CFD. By reducing expertise barriers, ChatCFD enhances CFD accessibility and provides methodological insights for AI-driven engineering simulations.
github: The code for ChatCFD is available at: https://github.com/EarlFan/ChatCFD.
Keywords large language model, multi-agent system, computational fluid dynamics, automated CFD
1 Introduction
Computational Fluid Dynamics (CFD) is a cornerstone technology in diverse scientific and engineering disciplines, including aerospace [28, 14, 12], energy systems [17, 2], urban environment [22, 19, 27], combustion [26, 10, 11], and biomedical applications [24, 9, 7]. It provides indispensable tools for simulating intricate fluid behaviors, thereby enabling design innovation and scientific discovery [3, 4]. However, the practical deployment of CFD faces significant challenges. Traditional CFD simulations require extensive domain-specific expertise [32] and frequently rely on costly commercial software packages. Even experienced engineers must invest considerable time in critical tasks, including solver selection, model development, mesh generation, boundary condition specification, and post-processing[20]. These complexities, compounded by the substantial computational resources needed, restrict the broader adoption of CFD, particularly for smaller enterprises and research groups, thus creating high entry barriers and potentially stifling innovation [4]. Consequently, there is a critical need for automated, user-friendly, and cost-effective CFD tools. To address this, we introduce ChatCFD, an automated agent system that leverages advanced artificial intelligence to streamline the CFD workflow.
Recent advances in Artificial Intelligence (AI) have revolutionized the automation landscape of complex scientific workflows. At the forefront of this transformation are Large Language Models (LLMs)—including GPT [1], Gemini [30], and DeepSeek [13]—alongside advanced multi-agent frameworks such as MetaGPT [15], and AutoGen [34]. These LLMs exhibit remarkable proficiency in natural language understanding, code generation, and complex reasoning tasks through chain-of-thought processes [33], enabling the development of autonomous LLM-based agents that can effectively perceive, reason, and act within their operational environment. Furthermore, the integration of Retrieval-Augmented Generation (RAG) [21] significantly enhances these agents. RAG is a technique where, before generating a response or performing an action, the LLM retrieves relevant information from an external, often domain-specific, knowledge base. This retrieved context is then provided to the LLM to inform its output. This process strengthens agents by grounding their responses in factual data, reducing hallucinations, and enabling them to perform more effectively in specialized domains like CFD.
This technological convergence has catalyzed the development of powerful LLM-based agents capable of automating the CFD workflow spectrum, from interpreting natural language specifications to configuring and executing simulations. This advancement represents a significant step toward making complex CFD simulations more accessible and efficient. OpenFOAM [18], in particular, stands out as the most popular open-source and free CFD software package, whose extensive capabilities, flexible architecture, and vibrant community support have made it a preferred choice for academic research, industrial applications, and consequently, a common framework for developing novel CFD automation agents. This accessibility has fostered innovation, leading to several pioneering initiatives exploring the integration of LLMs within the CFD domain. For example, MetaOpenFOAM [5] was introduced as an LLM-based multi-agent framework utilizing the MetaGPT [15] paradigm and RAG to automate CFD simulation tasks from natural language inputs, primarily demonstrated with OpenFOAM tutorial cases. Concurrently, OpenFOAMGPT [25] explored RAG-augmented LLM agents for OpenFOAM-centric CFD simulations, assessing model proficiency in tasks like zero-shot case setup and boundary condition adjustments. Foam-Agent [23] further advanced this area by proposing a multi-agent system with innovations such as hierarchical multi-index retrieval and dependency-aware file generation. Other research avenues have concentrated on fine-tuning LLMs with domain-specific datasets, like NL2FOAM [8], to enhance direct translation from natural language to executable CFD configurations, potentially using techniques like LoRA [16]. Despite these significant strides in automating CFD pipeline components and reducing technical barriers, existing systems often struggle with the end-to-end management of complex simulations without human intervention.
Existing CFD automation agents are often limited by their focus on rudimentary tasks, typically restricted to OpenFOAM tutorial-level examples, raising concerns about their generalization to complex, unseen cases. Three critical challenges remain unaddressed. First, describing CFD cases using brief natural language is often inadequate, as practical simulations require precise specifications beyond simplified verbal descriptions. If not given a detailed and professional task description, effective agents must interactively clarify user intent to define simulation requirements accurately. Second, handling intricate geometries poses significant difficulties, as mesh generation for real-world scenarios exceeds the capabilities of basic OpenFOAM utilities like “blockMesh” and demands specialized meshing tools. The inability to integrate user-defined or externally generated meshes limits these agents’ utility for diverse, practical CFD problems. Third, incorporating domain-specific knowledge into CFD agents remains challenging. Engineering-scale simulations require comprehensive initial and boundary condition specifications, necessitating domain-specific structural thinking and seamless integration of specialized domain expertise into the automation pipeline. These shortcomings highlight a critical gap in current CFD automation frameworks.
Figure 1: Overview of the ChatCFD automated agent system for streamlining CFD simulations within the OpenFOAM framework. ChatCFD enables researchers and engineers to configure and execute simulations with minimal CFD or OpenFOAM expertise. The system comprises three core components: (1) an interactive chat interface for users to input case descriptions or upload mesh files, (2) a thinking system, the core decision-making module (detailed in Figure 2), and (3) a simulation engine that executes cases, collects error logs, and delivers final results.
To address the limitations outlined above, this paper introduces ChatCFD, an automated agent system for configuring and executing CFD simulations within the OpenFOAM framework, as illustrated in Fig. 1. ChatCFD enables users to define cases interactively via a chat interface or by uploading detailed technical files, such as well-documented research papers. Leveraging large language models (e.g., DeepSeek-R1, DeepSeek-V3), a sophisticated multi-agent architecture, and specialized OpenFOAM knowledge, ChatCFD autonomously manages the entire simulation workflow. It adeptly handles complex CFD setups through iterative trial-and-error, supporting comprehensive initial and boundary conditions, intricate physical models, and user-defined or externally generated mesh files. This architecture enables ChatCFD to address diverse, unseen scenarios and adapt various physical models, marking a significant advancement over prior agents.
This paper presents the detailed architecture of the ChatCFD system and validates its modular effectiveness through experimental analysis. The paper is organized as follows: Section 2 details the system architecture, Section 3 provides experimental results demonstrating the effectiveness of its components, and Section 4 summarizes the key findings. An appendix includes illustrative examples and prompts to clarify ChatCFD’s operational functionality.
2 Method
Figure 2: Architecture of the ChatCFD framework for automated CFD simulations, illustrating the four-stage workflow and agent structure. The stages are: (1) Knowledge Base Construction, creating a JSON database from OpenFOAM manuals and tutorials; (2) User Input Processing, enabling user interaction via natural language or document and mesh uploads; (3) Case File Initialization, generating OpenFOAM case files using the knowledge base; and (4) Simulation Execution and Error Reflection, running simulations, converting meshes with fluentMeshToFoam, and resolving errors (dimension mismatches, missing files, persistent errors, general issues) using RAG-based modules ReferenceRetriever and ContextRetriever. The agent structure integrates DeepSeek-R1 and DeepSeek-V3 for intelligent processing, with iterative error correction.
ChatCFD is an automated CFD agent system that leverages the OpenFOAM framework to process multi-modal user inputs, including research articles and mesh files, to configure and execute CFD simulations based on user instructions. As illustrated in Fig. 2, the ChatCFD framework implements a comprehensive workflow to streamline simulation setup, execution, and analysis.
2.1 Stage 0: Knowledge Base Construction
The preprocessing aims to establish a foundational knowledge base for CFD tasks, performed in advance to optimize ChatCFD’s operations. Ideally, this knowledge base contains general CFD principles (e.g., numerical differentiation, discretization schemes), OpenFOAM-specific manuals, and a comprehensive example library. Fine-tuning large language models (LLMs) for general CFD knowledge is beyond the current scope; readers are referred to recent work by Dong et al. [8]. This study focuses on preprocessing OpenFOAM manuals and publicly accessible examples. A systematic analysis of OpenFOAM manuals and tutorials constructs a structured JSON database, cataloging essential parameters such as solver configurations, turbulence models, and file dependencies, providing a critical foundation for accurate case setup and error correction.
Initially, preprocessed manual and tutorial data define the file dependency relationship, accessed during the RAG process to enhance the agent’s error correction ability. To ensure robust operation, strategic case filtering is implemented, removing OpenFOAM-specific headers, excluding cases with external dependencies or auxiliary folders to maintain independence, and filtering out cases with extensive non-uniform field definitions that reduce LLM response quality without contributing to configuration. Filtering criteria and implementation details are provided in Appendix A.
Subsequently, the example library is preprocessed through systematic extraction and organization. All case configuration files are parsed and converted into a structured JSON format, tagged with metadata such as solver specifications, turbulence model types, and flow regime characteristics. Analysis of solver and turbulence model distributions determines specific file requirements for each configuration. For instance, the simpleFoam solver requires configuration files including system/fvSolution, system/controlDict, system/fvSchemes, 0/p, 0/U, constant/transportProperties, and constant/turbulenceProperties. Similarly, the kOmegaSST turbulence model requires system/fvSolution, system/controlDict, system/fvSchemes, 0/p, 0/U, 0/k, 0/omega, 0/nut, constant/turbulenceProperties, and constant/transportProperties.
2.2 Stage 1: User Input Processing
The initial stage of ChatCFD features an interactive, multi-modal interface that enables users to define CFD simulations by either conversing with the DeepSeek-R1 model or uploading documents and mesh files, built using the Streamlit Python framework. This interface leverages the preprocessed knowledge base to facilitate accurate case specification with minimal CFD expertise. The knowledge base, comprising OpenFOAM manuals, tutorial cases, and a structured JSON database of solver configurations and turbulence models, informs the system’s natural language processing and case extraction, ensuring robust and context-aware interactions. The workflow comprises four key phases:
  • Input Submission: Users can upload research articles or describe the case through dialogue to provide the basis for case analysis.
  • Case Extraction: The system extracts and catalogs CFD cases from uploaded documents or conversation inputs, identifying solver configurations and physical models, and presents them in a structured format with unique identifiers (e.g., Case 1, Case 2).
  • Case Selection: Through natural language interaction, the system guides users to select a target case, provides detailed specifications for verification, and confirms the selection.
  • Mesh Integration: The system assists users in uploading mesh files, currently supporting Fluent-format .msh files.
Upon completing this workflow, ChatCFD aggregates three essential components for downstream processing, as shown in Fig. 2:
  • Case Documentation: The system retains the full research article or conversation details, providing access to all technical specifications, including solvers, boundary conditions, and simulation parameters.
  • Case Specification: The selected case is recorded with detailed metadata, including solver types, numerical schemes, physical models, and source references, ensuring precise setup.
  • Mesh Data: Robust file transfer and path management ensure reliable delivery of mesh data to subsequent stages.
2.3 Stage 2: Case File Initialization
The CFD engineer layer, powered by the DeepSeek-R1 model, initializes OpenFOAM case files through a streamlined two-phase process, leveraging the preprocessed knowledge base.
In the first phase, the system generates a configuration list of required case files by analyzing the case description from Stage 1 to determine the appropriate solver and turbulence model. It queries the knowledge base’s OpenFOAM dataset to compile a tailored list of necessary files (e.g., system/fvSolution, 0/p) for the solver-turbulence model pair, enhancing initialization accuracy and minimizing missing file errors. Such a procedure, based on an explicit manner, substantially enhances the quality of case file initialization and reduces the incidence of missing file errors.
The second phase extracts detailed case setups from three sources: the configuration list, user prompt, and example library. The configuration list ensures precise case identification, while the prompt is segmented by section and filtered using the sentence-transformers/all-mpnet-base-v2 library with relevance thresholds. Hierarchical parameter extraction follows three steps:
  • Boundary Condition Validation: The system identifies boundary types from the prompt, verifies compatibility with OpenFOAM v2406, and enforces strict naming conventions to avoid errors. For example, in the airfoil case shown in Fig. 3, while the LLM may default to using a generic “inlet” type for the Inlet boundary based on common practice, the system ensures it correctly implements the specified “freeStream” boundary condition type to maintain fidelity with the original case setup.
  • Field Value Specification: Initial and boundary condition values are extracted for all field files (e.g., 0/p, 0/nut, 0/nuTilda) listed in the configuration.
  • Configuration File Generation: All required OpenFOAM files are generated, including field files (0/), physical properties (constant/, excluding polyMesh/), and control parameters (system/*).
Step 3 leverages DeepSeek-R1’s reasoning capabilities to orchestrate the simultaneous generation of all required case configuration files while ensuring consistency across physical models and boundary conditions. For numerical configurations not explicitly specified in the research article, such as discretization schemes and solution algorithms defined in system/fvSchemes and system/fvSolution files, ChatCFD employs DeepSeek-R1’s advanced reasoning capabilities to construct appropriate configurations based on its understanding of CFD best practices and the physical requirements of the case. The agent maintains coherence between interdependent parameters while adapting numerical schemes to match the flow physics and solution requirements indicated by the article’s methodology.
2.4 Stage 3: Simulation Execution and Error Reflection
Stage 3 executes the CFD simulation and analyzes runtime logs for automated error correction through three sequential steps:
  • Case Deployment: Case files from Stage 2 are deployed to the designated directory, and uploaded mesh files are converted to OpenFOAM format using fluentMeshToFoam, storing results in constant/polyMesh.
  • Temporal Configuration: Temporal parameters in system/controlDict are set, running 10 iterations/time steps for steady-state or transient simulations, with outputs at steps 0, 5, and 10. Transient cases enforce a maximum CFL number of 0.6, with time steps of $1\times 10^{-5}$ s for incompressible flows and $1\times 10^{-8}$ s for compressible flows.
  • Simulation Execution: The simulation runs using the specified OpenFOAM solver, with execution output logged to case_run.log for analysis.
ChatCFD automates error detection and correction in OpenFOAM case files by analyzing runtime logs using RAG and two retrieval modules:
The ReferenceRetriever module queries the preprocessed OpenFOAM knowledge database to select reference case files with matching or similar solvers and physical configurations. Typically, three reference files balance guidance and prompt complexity, ensuring contextually relevant corrections (e.g., distinguishing incompressible vs. compressible flows).
The ContextRetriever module compiles all current case configurations into a structured JSON format, providing the DeepSeek-V3 model with comprehensive context to address errors like missing files, especially in physically coupled problems (Section 3.3).
Upon detecting a runtime error, ChatCFD classifies it into four categories, detailed in Appendix B with examples:
  • Dimension Mismatches: Inconsistent physical dimensions in field or configuration files are identified by DeepSeek-V3 through log pattern analysis. Corrections involve adjusting dimensions to match reference files from ReferenceRetriever, ensuring consistency.
  • Missing Files: Missing case files, as reported in MetaOpenFOAM [5], are detected via log messages. DeepSeek-V3 identifies the missing file, ReferenceRetriever provides a template, and ContextRetriever supplies existing case details. DeepSeek-R1 generates the file, ensuring consistency in boundary patches, dimensions, and solver settings.
  • Persistent Errors: Errors recurring across correction cycles are tracked by DeepSeek-R1. The problematic file is identified, and DeepSeek-V3 rewrites it using reference files from ReferenceRetriever and the original file content.
  • General Errors: Errors like incorrect keywords or numerical schemes, comprising  50% of runtime issues, are addressed through a three-step process:
    • Error Localization: DeepSeek-R1 analyzes the log and case files (via ContextRetriever) to pinpoint the erroneous file, constraining the search to relevant files to reduce LLM errors.
    • Correction Proposal: DeepSeek-R1 generates step-by-step correction advice using the error log, erroneous file, and reference files from ReferenceRetriever, preserving initial and boundary conditions from Stage 1.
    • Correction Application: DeepSeek-V3 applies the correction, adhering to the advice and maintaining dimensional consistency.
2.5 LLM Configurations
ChatCFD employs a dual-model approach, integrating DeepSeek-R1 and DeepSeek-V3 to handle the initiation and correction of case files in the workflow. The application of these two LLMs follows distinct principles. DeepSeek-R1 takes the lead in Stage 2, where it extracts comprehensive case setups from textual descriptions and configures the initial case files. Its capacity for in-depth text comprehension and error correction is also crucial in Stage 3, where it analyzes and rectifies general errors and generates any missing files. DeepSeek-V3’s contributions are concentrated in Stage 3, focusing on the correction of case files. It excels at managing highly patterned tasks, such as identifying dimensional and file-missing errors. Furthermore, DeepSeek-V3 handles specific corrections, such as adjusting field file dimensions, rebuilding case files to correct persistent errors, and amending files according to precise step-by-step instructions.
This synergistic use of DeepSeek-R1 and DeepSeek-V3 significantly enhances ChatCFD’s overall performance, particularly in correcting case files. While DeepSeek-R1 excels in Stage 3 with its strong reasoning and planning for long-text comprehension and comprehensive file configuration, its tendency to hallucinate or generate inaccuracies became a hurdle during development. These hallucinations led to problems such as misidentifying dimensional errors and inserting unwanted markdown formatting into corrected files, issues that extensive prompt engineering alone could not resolve.
To address these challenges, ChatCFD adopted a strategic hybrid method. In the case file correction module, DeepSeek-V3 now manages the majority of error corrections, which has markedly increased the success rate of these corrections. DeepSeek-R1 retains responsibility for resolving file-missing errors, as this task requires a holistic analysis of all configuration files for consistency—a strength of DeepSeek-R1’s content processing. A similar division applies to error type identification: DeepSeek-V3 handles pattern-based errors like dimension and file-missing issues, while DeepSeek-R1 tackles complex errors, such as repetitive and common problems, which necessitate broader content analysis.
For DeepSeek-V3, the thought prompting by Wu et al. [35] in 2024 is used to equip DeepSeek-V3 with some thinking abilities. An auxiliary function then helps to extract the final answer from DeepSeek-V3’s response. The thought prompting text instructs the model to respond comprehensively and in detail, and to first write down its thought process after the phrase "Here is my thought process:" before providing its actual response after "Here is my response:". An example of thought prompting is given in Appendix C.
Furthermore, ChatCFD maintains a log of all Question-Answer (QA) histories with the LLMs for each CFD case. These QA logs are valuable for analyzing the LLM’s behavior as it configures CFD cases within the OpenFOAM framework.
The experimental validation detailed in Section 3 was conducted on the VolcEngine platform [31] from March 1 to March 12, 2025. The pricing structure for the models on this platform, used for evaluating token usage and cost, is as follows: for DeepSeek-R1, the cost is \$0.00055 (0.004 RMB) per 1,000 input tokens and \$0.0022 (0.016 RMB) per 1,000 output tokens; for DeepSeek-V3, the cost is \$0.00021 (0.0015 RMB) per 1,000 input tokens and \$0.00082 (0.006 RMB) per 1,000 output tokens. All token usage and cost calculations mentioned in Section 3 are based on this pricing structure from the VolcEngine platform.
3 Results and discussion
The performance of ChatCFD was assessed through validation experiments on basic CFD examples and two complex cases, using published literature as prompts. ChatCFD achieved error-free configurations in 30-40% of cases and operational success in 60-80%, marking a significant advance in automated CFD setup. The system demonstrated robust handling of complex configurations, particularly those with diverse turbulence models, highlighting its adaptability to varied physical scenarios. Analysis revealed two key performance factors: the extent of physical coupling in the flow and the availability of comprehensive knowledge about relevant physical models. These findings elucidate the capabilities and limitations of AI-driven CFD automation, contributing to the advancement of automated CFD workflows.
3.1 CFD cases for Experiments
Figure 3: Demonstration of the two tested CFD cases, including grids of (a) the Naca0012 case [29] and (b) the Nozzle case [36], and boundary types of (c) the Naca0012 case and (d) the Nozzle case.
We selected two representative test cases for simulation: an incompressible flow and a compressible flow case (illustrated in Figure 3). These cases were chosen based on the comprehensive documentation available in their source articles, which provided detailed specifications of the solver configurations, turbulence models, and mesh characteristics essential for accurate CFD reproduction. For the incompressible flow case, we utilized the work of Sun et al.[29], who conducted simulations of a NACA0012 airfoil at 10° angle of attack employing simpleFoam solver with the SpalartAllmaras turbulence model. For the compressible flow case, we referenced Yu et al.[36], who performed simulations of a nozzle with pressure ratio 3. The rhoCentralFoam and SpalartAllmaras model are used in the Nozzle case.
Given the documented challenges in modifying OpenFOAM turbulence models using LLM, as reported by Pandey et al.[25], we extended our investigation to include additional experiments with the NACA0012 case implementing various turbulence models. Table 1 presents a comprehensive overview of all case configurations and experimental iterations conducted in this study. High-quality computational meshes for these experiments were generated using commercial software. The experimental design featured a structured testing protocol: Cases 1 and 5 were subjected to an extensive model ablation analysis, with 10 iterations performed for each of the five system configurations. In contrast, Cases 2 through 4 were evaluated using only the complete system configuration, also with 10 iterations each. This systematic methodology yielded a total of 130 ChatCFD experimental runs, forming a robust dataset for subsequent analysis.
Table 1: Summary of Test Cases for ChatCFD Validation
Case ID Case Mesh Physical Models Experimental Design
Case 1 NACA0012
simpleFoam with
SpalartAllmaras
10 runs $\times$ 5 system
configurations
Case 2 NACA0012
simpleFoam with
kOmegaSST
10 runs $\times$ the complete
system configuration
Case 3 NACA0012
simpleFoam with
kEpsilon
10 runs $\times$ the complete
system configuration
Case 4 NACA0012
simpleFoam with
RNGkEpsilon
10 runs $\times$ the complete
system configuration
Case 5 Nozzle
rhoCentralFoam with
hePsiThermo and
SpalartAllmaras
10 runs $\times$ 5 system
configurations
3.2 Model Ablation Analysis
To systematically evaluate the contributions of different components and strategies, ablation experiments were performed using five distinct system configurations. These configurations vary in terms of comprehensive paper interpretation detailed in Stage 2 (Initiating Case Files), the sophistication of error handling logic, and the utilization of retrieval modules (ReferenceRetriever and ContextRetriever) in Stage 3 (Correcting Case Files):
  • Configuration A (Baseline System): Utilizes a simplified parameter extraction process instead of the three-step process in Stage 2. In Stage 3, error correction is limited to a basic “general error” pathway, with both ReferenceRetriever (tutorial-based retrieval) and ContextRetriever (current case retrieval) modules deactivated.
  • Configuration B (ReferenceRetriever with Basic Error Handling): Maintains the simplified parameter extraction (Stage 2) and basic ’general error’ handling (Stage 3) of Configuration A, but activates the ReferenceRetriever module to inform corrections with tutorial case data.
  • Configuration C (Multi-Category Errors, ReferenceRetriever, Simplified Extraction): Employs the simplified parameter extraction in Stage 2. However, Stage 3 incorporates the full multi-category error correction logic (addressing dimensional, missing file, repetitive, and general errors) and activates the ReferenceRetriever module. The ContextRetriever module remains deactivated.
  • Configuration D (Multi-Category Errors, ReferenceRetriever & ContextRetriever, Simplified Extraction): Also uses the simplified parameter extraction in Stage 2. Stage 3 employs the full multi-category error correction logic, supported by both ReferenceRetriever and ContextRetriever modules.
  • Configuration E (Complete System): Represents the most comprehensive setup. It features comprehensive paper interpretation in Stage 2. In Stage 3, it utilizes the full multi-category error correction logic, augmented by both ReferenceRetriever and ContextRetriever modules.
Figure 4: Success rates of ChatCFD in reproducing NACA0012 Case 1 and Nozzle Case 5 across five distinct Configurations A to E. The horizontal axis delineates these configurations, while the vertical axis quantifies the success rates. A 10-step operation indicates the CFD case successfully advanced 10 time steps or executed 10 iteration steps. Accuracy denotes precise case configuration in accordance with the source article’s specifications.
Figure 4 illustrates the performance of ChatCFD across the five distinct system Configurations A through E for two benchmark test cases: NACA0012 Case 1 and Nozzle Case 5. Two primary metrics were employed for evaluation: the 10-step operational success rate, signifying the capability to complete ten simulation steps (time steps for transient cases or iterations for steady-state cases) without critical errors, and the accurate configuration success rate, indicating that the CFD case was set up in exact accordance with the specifications detailed in the source research article. The 10-step criterion was adopted to enable rapid assessment of operational stability for the system’s reflection mechanism, as target simulation durations are often unspecified in the literature. It was observed that configurations meeting this 10-step milestone were generally free of major operational faults, though minor discrepancies in boundary conditions parameters from the article’s description could still occur. For cases achieving accurate configuration, the simulation results demonstrated good agreement with published data, as detailed in Appendix D.
For the NACA0012 Case 1, the data in Figure 4 show a clear progression in the 10-step operational success rate with increasing system sophistication. The baseline Configuration A achieved a 20% success rate. This rate systematically improved to 40% for Configuration B, 60% for Configuration C, and reached 80% for Configuration D, which benefited from enhanced retrieval capabilities. This trend highlights the positive impact of incrementally integrating retrieval modules and OpenFOAM-specific knowledge. However, despite its high operational success, Configuration D failed to achieve any accurate configuration success, being unable to precisely replicate the case setup from the source article. This crucial gap was addressed by Configuration E, which incorporates comprehensive article interpretation and parameter extraction as part of its advanced Stage 2 processing. With Configuration E, the accurate configuration success rate for the NACA0012 case improved significantly to 40%, while maintaining the 80% operational success rate.
The comprehensive article interpretation module in Stage 2 is integral to achieving accurate CFD configurations, particularly for cases involving non-standard boundary conditions. For instance, the NACA0012 case, as depicted in Figure 3, employs several specialized OpenFOAM boundary conditions (freestream, freestreamVelocity, and freestreamPressure) that deviate from common practices. Analysis of the LLM’s internal reasoning processes indicated that, without robust article interpretation capabilities and strict adherence to documented specifications, the model tends to default to more conventional boundary conditions. This can lead to erroneous configurations, such as using standard inlet-type conditions for inlet velocities and pressureOutlet-type conditions for outlet pressures, instead of the requisite specialized freestream conditions.
This tendency for conventional settings also extends to numerical parameters, including inlet velocity values. To mitigate these issues, the comprehensive article interpretation module was developed in Stage 2, incorporating a three-step process for analyzing articles. This process emphasizes the precise extraction of case specifications and ensures adherence to OpenFOAM’s inherent boundary condition types, as detailed in Section 2. The introduction of this module yielded substantial improvements: while the 80% success rate for the 10-step operational benchmark was maintained, the system’s capability to accurately replicate documented case configurations increased dramatically from 0% to 40%. This significant enhancement underscores the efficacy of the refined interpretation strategy in correctly identifying and implementing specialized case requirements.
In the Nozzle Case 5, as depicted in Figure 4, the introduction of the ContextRetriever module led to a substantial improvement in the 10-step operational success rate. Configurations lacking ContextRetriever (i.e., up to Configuration C) exhibited a 0% success rate for this metric. The integration of ContextRetriever in Configuration D elevated this rate to 60%. This marked improvement is attributed to ContextRetriever’s enhanced capability to diagnose and address complex coupling-related errors (refer to Appendix B for examples), a notable limitation in earlier configurations. For instance, when encountering thermophysical parameter errors in OpenFOAM (such as those related to $\Psi$, the compressibility), previous configurations often misidentified the error source as the constant/thermodynamicProperties file. However, such errors frequently originate from inconsistencies in the equation of state specifications within boundary condition entries in field files (e.g., 0/p or 0/T). The ContextRetriever module overcomes this by performing a comprehensive analysis across all relevant case directories (0/, constant/, and system/), thereby enabling more accurate error localization and effective correction strategies.
The differential performance of the ContextRetriever module when applied to the NACA0012 Case 1 and the Nozzle Case 5 can be attributed to fundamental distinctions in their underlying flow physics. NACA0012 Case 1, an incompressible flow simulation employing the pisoFoam solver, exhibits a comparatively weak coupling between pressure and velocity fields. In contrast, Nozzle Case 5, a compressible flow scenario utilizing the rhoCentralFoam solver, is characterized by strong interdependencies among density, temperature, and pressure, governed by the equation of state. The advanced analytical capabilities of ContextRetriever, particularly in discerning inter-file dependencies and identifying errors related to coupled physics, proved notably more effective for the Nozzle case due to its inherent complexities. While still beneficial, the module’s impact was less pronounced for the NACA0012 case, which features a less coupled system. This observed variance underscores the proficiency of ContextRetriever in managing comprehensive physical information across a spectrum of problem complexities.
Advancing from Configuration D to E for the Nozzle Case 5 preserved the 60% 10-step operational success rate while notably enhancing the accurate configuration success rate from 0% to 30%. This pattern of improvement in configuration precision, achieved by incorporating comprehensive article interpretation in Stage 2, mirrors the observations from the NACA0012 case, indicating a consistent benefit of the complete system configuration across different flow regimes.
Figure 5: Results for NACA0012 Case 1 using Configurations A to E, showing averages of 10-step operational success runs. Metrics shown include: number of reflection iterations (#reflection), total token consumption for DeepSeek-R1 and DeepSeek-V3 models (R1/V3 total token), execution cost (Cost \$), and average token consumption per reflection round for both models (R1/V3 token/#reflection)
Figure 5 presents a detailed analysis of the NACA0012 Case 1 performance across the five distinct Configurations (A through E), based on averaged results from 10-step operational success runs. This analysis yields several key insights into the system’s operational efficiency and cost-effectiveness.
The evolution of reflection iterations, illustrated in Figure 5(a), demonstrates an encouraging improvement pattern. As retrieve modules were integrated from configuration A to C, we observed a substantial decrease in required reflection iterations, followed by stabilization. This observation aligns with MetaOpenFOAM’s findings [6] regarding RAG positive impact on LLM response quality and CFD agent performance. The subsequent stability in reflection counts from configurations C to E suggests that additional RAG implementations, while beneficial for other aspects, had minimal impact on ChatCFD’s reflection behavior. This plateau is reasonable given that ContextRetriever and comprehensive paper interpretation modules were primarily designed to enhance performance on physical-coupled cases and setup accuracy respectively, rather than affecting reflection behaviors for weak coupling cases.
An analysis of token consumption and associated costs, depicted in Figure 5(b), reveals a pattern of improving resource utilization efficiency. Transitioning from Configuration A to D, while the token consumption of the DeepSeek-R1 model remained relatively stable, the DeepSeek-V3 model’s consumption decreased markedly. This reduction contributed to lower per-case operational costs (decreasing from \$0.078 to \$0.071). This cost improvement, concurrent with a significant increase in operational success rates from 20% to 80%, underscores the enhanced overall effectiveness of the ChatCFD system.
The transition from Configuration D to E resulted in a notable enhancement of system accuracy, as illustrated in Figure 4. This advancement was primarily achieved by improving the paper interpretation module’s efficacy in extracting CFD case setups during Stage 2 and by ensuring greater consistency of these setups throughout the file correction processes in Stage 3. Although this upgrade led to increased DeepSeek-R1 token consumption and a corresponding 30% rise in operational cost (from \$0.071 to \$0.091), the substantial improvement in accurate case configuration—from 0% to 40%—justifies this investment.
Figure 5(c) illustrates the token consumption patterns per reflection iteration for both DeepSeek-R1 and DeepSeek-V3 models across the different Configurations. From Configuration A to E, the DeepSeek-R1’s consumption increases to more than two times, where the increasing consumption are mostly due to the integration of the ContextRetriever module in Configuration C. The DeepSeek-V3’s consumption initially rose due to expanded error handling capabilities and reference case integration through ReferenceRetriever. The subsequent stabilization of DeepSeek-V3 token usage from Configuration C to E reflects the maturation of the reflection module’s design, achieving optimal operational efficiency without necessitating additional DeepSeek-V3 module invocations.
Figure 6: Results for Nozzle Case 5 using Configurations D and E, showing averages of 10-step operational success runs. Metrics shown include: number of reflection iterations (#reflection), total token consumption for DeepSeek-R1 and DeepSeek-V3 models (R1/V3 total token), execution cost (Cost \$), and average token consumption per reflection round for both models (R1/V3 token/#reflection)
Figure 6 presents a comparative performance analysis for the Nozzle Case 5, specifically contrasting Configurations D and E, based on simulations that successfully completed ten steps. Configurations A through C were omitted from this comparison due to their inability to consistently reach this ten-step benchmark. The key differentiator between Configurations D and E is the integration of the comprehensive paper interpretation module’s enhanced article interpretation capabilities in Configuration E. Notwithstanding this enhancement, Figure 6(a) indicates that the number of reflection iterations remained largely comparable between these two configurations, with only marginal reductions observed for Configuration E.
The deployment of comprehensive paper interpretation module resulted in increased token utilization, as depicted in Figure 6(b). Configuration E demonstrated higher token consumption by the DeepSeek-R1 model, whereas the DeepSeek-V3 model’s token usage remained stable across both Configurations D and E. In terms of operational cost, the average per-case expenditure saw a modest increase from \$0.28 to \$0.31, an approximate 10% increment. Nevertheless, this additional investment yielded significant improvements in performance, particularly in case setup accuracy, which, as shown in Figure 4, rose substantially from 0% to 30%.
Further examination, illustrated in Figure 6(c), reveals that the average token consumption per reflection iteration for both the DeepSeek-R1 and DeepSeek-V3 models was comparable across Configurations D and E. This consistency underscores the efficacy of the system’s architectural design, indicating that the comprehensive paper interpretation module detailed interpretation component operates as a distinct layer that augments overall system performance without adversely affecting the fundamental reflection mechanisms.
3.3 Performance degradation in configuring physical-coupled CFD cases
Table 2: Statistical averages of accurate configured cases for Naca0012 Case 1 and Nozzle Case 5.
Metrics Naca0012, Case 1 Nozzle, Case 5
Physical models Flow type: Incompressible Solver: simpleFoam Turbulence: SpalartAllmaras Flow type: Compressible Solver: rhoCentralFoam Turbulence: SpalartAllmaras Thermo model: hePsiThermo
Number of configuration files 9 12
Average tokens for all configuration files per case 1,792 2,647
Configuration files 0/p 0/U 0/nut 0/nuTilda constant/transportProperties constant/turbulenceProperties system/controlDict system/fvSchemes system/fvSolution 0/p 0/U 0/nut 0/nuTilda 0/T 0/alphat constant/transportProperties constant/turbulenceProperties constant/thermodynamicProperties system/controlDict system/fvSchemes system/fvSolution
DeepSeek-R1 call count 19.1 48.3
DeepSeek-V3 call count 17.4 35.2
Figure 7: ChatCFD’s average performance per execution when accurately configuring NACA0012 Case 1 and Nozzle Case 5. Metrics shown include: number of reflection iterations (#reflection), total token consumption for DeepSeek-R1 and DeepSeek-V3 models (R1/V3 total token), execution cost (Cost \$), and average token consumption per reflection round for both models (R1/V3 token/#reflection).
The model ablation analysis in Section 3.2 highlights significant performance variations among configurations A through E when applied to two distinct test cases: the incompressible NACA0012 Case 1 and the compressible Nozzle Case 5. These discrepancies are attributed to fundamental differences in their underlying flow physics, with subsequent discussion detailing the influence of physical coupling—a key characteristic of compressible flow prominent in the Nozzle case—on ChatCFD’s performance.
The Nozzle case utilizes the rhoCentralFoam solver, which inherently manages strong interdependencies between pressure, temperature, and density through its thermophysical models. Its compressible flow solution also dictates supersonic or subsonic conditions based on local velocity and sound speed, which are coupled with thermodynamic properties (e.g., pressure, temperature). In contrast, the NACA0012 case employs the simpleFoam solver, featuring minimal such coupling. As shown in Table 2, these differing physical complexities translate to substantially different configuration requirements. The Nozzle case demands more configuration files with greater token content and three additional thermophysical-related files (0/T, 0/alphat, constant/thermodynamicProperties). Furthermore, achieving a precise configuration for this compressible flow scenario necessitates over twice the DeepSeek-R1 and DeepSeek-V3 model calls compared to the incompressible NACA0012 case.
Section 3.2 results demonstrate that the strong physical coupling in the Nozzle case correlates with diminished success rates. Figure 4 illustrates this, showing that the highly coupled Nozzle case consistently achieves lower 10-step success and precise configuration rates across ten experimental runs. This performance degradation is attributed to two primary factors: the increased complexity of managing a larger number of files with more extensive content, and the stringent requirement for maintaining consistency across interconnected thermophysical models and related parameters (e.g., in files such as 0/p, 0/T, and constant/thermodynamicProperties).
The greater physical complexity of the Nozzle case translates directly to substantially higher LLM consumptions than NACA0012 Case 1 (Figure 7). Specifically, the Nozzle simulation requires over double the reflection iterations for convergence, consumes approximately triple the total DeepSeek-R1 tokens, and uses six times the DeepSeek-V3 tokens. These escalated resource needs lead to the Nozzle case incurring roughly triple the operational costs. Such pronounced differences in resource utilization directly quantify the greater computational burden imposed by the strong physical coupling in compressible flow simulations relative to their incompressible counterparts.
Notably, the DeepSeek-V3 token consumption for the Nozzle case increased by a factor of approximately six, substantially exceeding the DeepSeek-R1 token increase of approximately threefold. Analysis of case execution QA histories reveals this disparity stems from a higher incidence of dimensional inconsistencies and persistent errors during Nozzle case configuration, which in turn necessitated more frequent invocations of the DeepSeek-V3 model. Figure 7(c) corroborates this, showing that the per-reflection DeepSeek-V3 token multiplier for the Nozzle case (about 2.7x) is obviously larger than its DeepSeek-R1 multiplier (about 2x) compared to the NACA0012 case.
The analysis of dimensional errors reveals a notably higher frequency in compressible flow scenarios, such as the Nozzle case, attributable to two fundamental factors. First, compressible flow physics imposes inherently more stringent requirements for achieving consistent setups; this involves not only the dimensional consistency and intricate coupling of primary fields like pressure, density, and velocity, but critically, their coherent integration with a thermodynamic model (such as an equation of state). This model, which interrelates these variables, often via temperature and specific energy, demands precise dimensional and physical alignment across all governing equations and material property definitions, a level of rigor that significantly elevates the complexity compared to typical incompressible simulations. Second, and significantly, the predominant focus on incompressible flow applications in typical CFD scenarios appears to have created an implicit bias in LLM training data. This bias manifests as a tendency for the LLM to default to incompressible flow configurations, even when handling compressible cases. A pertinent example of how this bias can degrade compressible flow configurations is the differing treatment of pressure dimensions: in OpenFOAM, incompressible flows often implement kinematic pressure $p_{k}$, which is dynamic pressure over density and has dimensions of $L^{2}T^{-2}$ whereas compressible flows use absolute dynamic (thermodynamic) pressure (p) with dimensions $ML^{-1}T^{-2}$. An LLM incorrectly applying the incompressible pressure dimension convention to a compressible case due to this training bias would introduce a fundamental dimensional error. Consequently, when LLMs encounter compressible flow scenarios, these embedded assumptions can lead to dimensional erros, exemplified by the pressure dimension mismatch, that require additional correction cycles to resolve.
3.4 Effects of turbulence models on ChatCFD’s performance
Figure 8: ChatCFD’s performance when applying four turbulence models to the NACA0012 Cases 1 to 4, where the results represent the average outcomes of cases precisely configured according to the paper specifications. The turbulence models include SpalartAllmaras, kEpsilon, kOmegaSST, and RNGkEpsilon. Metrics shown include: accuracy (success rate of precise configuration according to literature), number of reflection iterations (#reflection), total token consumption of the DeepSeek-R1 model (R1 total token), execution cost (Cost \$), token count of configuration files (#token of config files), and the number of cases incorporating each turbulence model in OpenFOAM tutorials (#Case of model in OF tutorial).
Figure 8 presents a detailed performance analysis of ChatCFD across four distinct turbulence models—Spalart-Allmaras (SA), kEpsilon, kOmegaSST, and RNGkEpsilon—applied to the NACA0012 benchmark (Cases 1 to 4). The evaluation focuses on simulations that achieved accurate configuration in accordance with published literature specifications, thereby excluding instances that only demonstrated rudimentary operational functionality (i.e., 10-step operational success). All experimental results reported herein were obtained using Configuration E, representing the complete agentic system.
As illustrated in Figure 8(a), the analysis reveals distinct success rates for accurate case configuration across the four turbulence models. The SA model achieved the highest success rate at 40%, followed by the kOmegaSST model (30%) and the kEpsilon model (20%). The RNGkEpsilon model exhibited markedly lower performance, with a success rate of 10%. Notably, while OpenFOAMGPT [25] has underscored the difficulties LLMs encounter when modifying turbulence models during case configuration, ChatCFD’s capability to successfully implement all four models attests to its efficacy in addressing these challenges, thereby validating the proposed architectural design and implementation methodology.
A detailed comparison of performance metrics, presented in Figure 8(a,b), delineates a clear behavioral divergence. The SA, kEpsilon, and kOmegaSST models displayed consistent performance profiles, characterized by an average reflection count of approximately 6, alongside comparable DeepSeek-R1 token utilization and associated computational expenditures. Conversely, the RNGkEpsilon model demanded substantially greater computational resources, evidenced by an average reflection count of 23.5—nearly fourfold higher than the other models—leading to correspondingly elevated token consumption and operational costs.
To elucidate the origins of these performance disparities, an analysis of underlying factors was conducted, with results presented in Figure 8(c). This investigation considered both the token counts of configuration files and the prevalence of each turbulence model within the OpenFOAM tutorial suite. Although configuration file sizes exhibited only modest variations—the SA model requiring around 1,800 tokens compared to approximately 2,000 for the others—this minor difference is insufficient to account for the substantial performance gap. A more significant factor emerged from the distribution of tutorial cases: whereas the SA, kEpsilon, and kOmegaSST models were each featured in 28 or more tutorial examples, the RNGkEpsilon model was represented in merely one instance.
The pronounced performance deficit observed for the RNGkEpsilon model is attributable to two principal factors. Firstly, its relative infrequency in practical engineering applications translates to a scarcity of representative examples in the LLM’s training dataset. This limitation impairs the LLM’s proficiency in configuring RNGkEpsilon cases effectively in a zero-shot manner (i.e., without RAG). Secondly, the model’s minimal presence within the OpenFOAM tutorials severely curtails the efficacy of the ReferenceRetriever module in sourcing appropriate reference configurations from this dataset. This is particularly detrimental for RNGkEpsilon-specific parameters, such as those found in the constant/turbulenceProperties file and associated numerical schemes within system/fvSchemes and system/fvOption. This dearth of tutorial-based guidance necessitates a greater dependence on the LLM’s intrinsic, yet less developed, understanding of the RNGkEpsilon model, culminating in diminished performance relative to more ubiquitously employed turbulence models.
4 Conclusion
This paper introduced ChatCFD, an innovative LLM-based CFD agent system designed to automate CFD simulations within the OpenFOAM framework. By integrating a multi-agent architecture, domain-specific knowledge bases, and advanced error correction mechanisms, ChatCFD successfully demonstrates the feasibility of end-to-end automation for both incompressible (NACA0012 airfoil) and compressible (supersonic nozzle) flow cases. Compared with antecedent CFD agent systems, such as MetaOpenFOAM [5] and OpenFOAMGPT [25], ChatCFD distinguishes itself through several key innovations:
Firstly, ChatCFD advances the application paradigm of LLM agents in engineering simulation through an interactive guidance mechanism that innovatively integrates multimodal inputs—PDF files, mesh files, and interactive dialogue—with underlying physical constraints. This mechanism, featuring a dedicated mesh input module and a reinforced context-memory dialogue system, offers a ChatGPT-like interface. This interface systematically guides users through essential stages, including literature analysis, case selection, the uploading of user-defined meshes, and the interactive specification of CFD model parameters, thereby addressing the complexities inherent in practical CFD applications. Consequently, this approach substantially reduces operational hurdles compared to traditional command-line driven methodologies.
Secondly, the system employs a collaborative framework of specialized AI agents that leverage deep CFD domain expertise. This strategy involves: (a) a comprehensive article interpretation module, utilizing an advanced reasoning LLM (e.g., DeepSeek-R1), for extracting CFD case parameters (e.g., boundary names, types, values) from academic papers, augmented by extensive OpenFOAM-specific boundary condition information to mitigate model hallucination, which is vital for precise CFD case replication; (b) preprocessing data from OpenFOAM tutorial cases to construct comprehensive feature libraries, thereby enriching the system’s understanding through RAG; and (c) a "global planning-local optimization" hierarchical framework. This framework strategically synergizes the advanced reasoning and text interpretation capabilities of one LLM with the high-fidelity, low-hallucination file modification strengths of another. For instance, when correcting configuration errors, DeepSeek-R1 first analyzes error messages and relevant files to pinpoint the erroneous file and suggest corrections, which DeepSeek-V3 then implements, ensuring both accuracy and reliability.
Thirdly, ChatCFD features a powerful multi-category error correction mechanism with two specialized RAG modules: ReferenceRetriever (accessing tutorial templates) and ContextRetriever (collecting current case configurations). This feature actively monitors simulation outputs in real-time, dynamically selecting appropriate and effective treatments by classifying typical case configuration problems. These include, but are not limited to, dimensional inconsistencies, missing file errors, and recurring errors indicative of deeper setup issues, thereby systematically addressing critical logical consistency challenges frequently encountered by LLMs in complex CFD computations.
Systematic ablation studies and case validations underscored the critical contributions of ChatCFD’s components. Comprehensive paper interpretation proved paramount for achieving high configurational accuracy, particularly for cases with rare or non-standard boundary conditions (e.g., “freestream” for NACA0012), improving accurate configuration success from 0% to 40% for the NACA0012 case and from 0% to 30% for the Nozzle case. The ContextRetriever module, which provides contextual awareness of current case files, was especially vital for diagnosing and rectifying complex, physically-coupled errors in compressible flow simulations; this boosted the Nozzle case’s 10-step operational success rate from 0% to 60%. Overall, the system achieved accurate configuration for approximately 30-40% of tested scenarios directly from literature, with higher operational success rates (e.g., 80% for NACA0012 with full system capabilities), establishing a valuable performance benchmark. Furthermore, the study quantified the cost-effectiveness of different system configurations. For instance, a 30% cost increase yielded a 40% improvement in accurate configuration for the NACA0012 case, justifying the investment for enhanced precision. These results highlight that while LLMs can automate aspects of CFD, their efficacy is significantly amplified when augmented with structured domain knowledge and robust error-handling mechanisms.
Key findings indicate that ChatCFD’s performance is significantly impacted by the complexity and representation of CFD cases in its knowledge base. Configuring physically-coupled compressible flow scenarios, like the Nozzle case, drastically degrades performance compared to incompressible ones due to more numerous and intricate configuration files and the critical need for consistency across interdependent thermophysical parameters. This results in lower success rates for accurate configurations, substantially increased LLM calls, higher token consumption (especially for error correction by DeepSeek-V3), more reflection iterations, and tripled operational costs, with dimensional errors exacerbated by the inherent rigor of compressible physics and an LLM training bias towards incompressible flow conventions. Similarly, performance varies markedly with the chosen turbulence model; those well-represented in OpenFOAM tutorials (e.g., Spalart-Allmaras with 40% success) yield better results and lower resource use. In contrast, models like RNGkEpsilon, scarce in both general LLM training data and specific tutorial datasets (only one example), exhibit poor success (10%) and demand significantly more computational resources, underscoring how limited data representation hampers ChatCFD’s configuration proficiency and RAG effectiveness.
ChatCFD pioneers computational automation, extending LLM applications to engineering challenges. By reducing expertise and software barriers, it democratizes access to advanced methodologies, fostering innovation. Its approaches—structured knowledge integration and context-aware RAG—offer insights for AI-driven engineering domains, such as simulations and optimizations. Future work includes expanding the knowledge base and developing standardized metrics to enhance automation, paving the way for AI-assisted scientific discovery.
ACKNOWLEDGEMENTS
This work is sponsored by the National Natural Science Foundation of China Grant No. 92470127, 92270203, the Open Project of the National Key Laboratory of Scramjet Technology No. WDZC6142703202403, and the Overseas Postdoctoral Talents Program in Guangdong.
AUTHOR DECLARATIONS
Conflict of Interest
The authors declare no conflicts of interest.
Author Contributions
E Fan: Methodology, Software, Data curation, Formal analysis, Resources, Writing - original draft.
Weizong Wang: Resources, Funding acquisition.
Tianhan Zhang: Conceptualization, Methodology, Formal analysis, Resources, Writing - Review & Editing, Visualization, Supervision, Funding acquisition
Data Availability
The data that support the findings of this study are available from the corresponding author upon reasonable request.

### v2

@@
-ChatCFD: an End-to-End CFD Agent with Domain-specific Structured Thinking
+ChatCFD: An LLM-Driven Agent for End-to-End CFD Automation with Domain-Specific Structured Reasoning
 Abstract
-Computational Fluid Dynamics (CFD) is critical for scientific and engineering advancements but is hindered by operational complexity and the need for extensive domain expertise.
-This paper presents ChatCFD, a large language model (LLM)-driven pipeline that automates computational fluid dynamics (CFD) workflows within the OpenFOAM framework, enabling users to configure and execute complex simulations from natural language prompts or published literature with minimal prior expertise.
-A key innovation is its ability to think structurally in database construction, configuration validation, and error reflection, which systematically integrates domain-specific CFD and OpenFOAM knowledge (e.g., solvers, turbulence models, file dependencies) with general LLMs, thereby enhancing accuracy and adaptability.
-More specifically, ChatCFD employs a four-stage workflow: (1) Knowledge Base Construction, creating a structured JSON database from OpenFOAM tutorials and manuals; (2) User Input Processing, guiding users via a multimodal interface (dialogue, documents, mesh files); (3) Case File Initialization, generating case files using a preprocessed knowledge base; and (4) Simulation Execution and Error Correction, running simulations and resolving errors with Retrieval-Augmented Generation (RAG).
-Powered by DeepSeek-R1 and DeepSeek-V3 models, a multi-agent architecture, and specialized OpenFOAM knowledge, ChatCFD features an interactive chat interface, hierarchical parameter extraction, and a robust error correction system addressing dimension mismatches, missing files, persistent errors, and general issues.
-Validation results demonstrate that ChatCFD autonomously reproduces published CFD literature results without human intervention, a challenging task involving complex, unseen configurations beyond basic CFD examples and general LLM capabilities.
-This achievement underscores the significant advancement enabled by its domain-specific structured thinking ability.
-In addition, validation experiments demonstrate error-free configurations in 30-40% of incompressible and compressible CFD cases and operational success in 60-80%, establishing a benchmark for automated CFD.
-By reducing expertise barriers, ChatCFD enhances CFD accessibility and provides methodological insights for AI-driven engineering simulations.
-github: The code for ChatCFD is available at: https://github.com/EarlFan/ChatCFD.
-Keywords large language model, multi-agent system, computational fluid dynamics, automated CFD
+Computational Fluid Dynamics (CFD) is essential for advancing scientific and engineering fields, but is hindered by operational complexity, high expertise requirements, and limited accessibility.
+This paper introduces ChatCFD, an automated agent system for OpenFOAM simulations that processes multi-modal inputs (e.g., research papers, meshes) via an interactive interface, leveraging DeepSeek-R1 and DeepSeek-V3 large language models (LLMs), a multi-agent architecture, and OpenFOAM knowledge.
+Its four-stage pipeline, Knowledge Base Construction, User Input Processing, Case File Generation, and Execution and Error Reflection, enables iterative trial-reflection-refinement for intricate setups, supporting diverse physical models and external meshes.
+Validation on 205 benchmark tutorial cases, 110 perturbed variants, and 2 literature-derived cases demonstrates ChatCFD’s 82.1% operational success rate on basic cases, significantly outperforming MetaOpenFOAM (6.2%) and Foam-Agent (42.3%), and 60–80% on literature-derived complex cases, with turbulence model studies showing a 40% success rate for common models vs.
+10% for rare ones like RNGk-$\epsilon$.
+Physics coupling analyses reveal higher resource demands for multi-physics-coupled cases, while LLM bias toward simpler setups introduces persistent errors, such as dimensional inconsistency.
+Ablation studies highlight the efficacy of RAG-based modules and reflection mechanisms.
+By automating hypothesis testing and parameter exploration, ChatCFD accelerates scientific discovery in fluid mechanics and engineering, addressing LLM limitations through structured design and demonstrating strong potential as a modular component in MCP-based agent networks for collaborative multi-agent systems, paving the way for scalable AI-driven CFD innovation.
+The code for ChatCFD is available at: https://github.com/ConMoo/ChatCFD.
+Keywords large language model, multi-agent system, computational fluid dynamics, OpenFOAM, automated CFD
 1 Introduction
-Computational Fluid Dynamics (CFD) is a cornerstone technology in diverse scientific and engineering disciplines, including aerospace [28, 14, 12], energy systems [17, 2], urban environment [22, 19, 27], combustion [26, 10, 11], and biomedical applications [24, 9, 7].
+Computational Fluid Dynamics (CFD) is a cornerstone technology in diverse scientific and engineering disciplines, including aerospace [1, 2], energy systems [3, 4], urban environment [5, 6, 7], combustion [8, 9], and biomedical applications [10, 11].
 It provides indispensable tools for simulating intricate fluid behaviors, thereby enabling design innovation and scientific discovery [12, 13].
-However, the practical deployment of CFD faces significant challenges.
-Traditional CFD simulations require extensive domain-specific expertise [32] and frequently rely on costly commercial software packages.
-Even experienced engineers must invest considerable time in critical tasks, including solver selection, model development, mesh generation, boundary condition specification, and post-processing[20].
-These complexities, compounded by the substantial computational resources needed, restrict the broader adoption of CFD, particularly for smaller enterprises and research groups, thus creating high entry barriers and potentially stifling innovation [4].
-Consequently, there is a critical need for automated, user-friendly, and cost-effective CFD tools.
-To address this, we introduce ChatCFD, an automated agent system that leverages advanced artificial intelligence to streamline the CFD workflow.
-Recent advances in Artificial Intelligence (AI) have revolutionized the automation landscape of complex scientific workflows.
-At the forefront of this transformation are Large Language Models (LLMs)—including GPT [1], Gemini [30], and DeepSeek [13]—alongside advanced multi-agent frameworks such as MetaGPT [15], and AutoGen [34].
-These LLMs exhibit remarkable proficiency in natural language understanding, code generation, and complex reasoning tasks through chain-of-thought processes [33], enabling the development of autonomous LLM-based agents that can effectively perceive, reason, and act within their operational environment.
-Furthermore, the integration of Retrieval-Augmented Generation (RAG) [21] significantly enhances these agents.
-RAG is a technique where, before generating a response or performing an action, the LLM retrieves relevant information from an external, often domain-specific, knowledge base.
-This retrieved context is then provided to the LLM to inform its output.
-This process strengthens agents by grounding their responses in factual data, reducing hallucinations, and enabling them to perform more effectively in specialized domains like CFD.
-This technological convergence has catalyzed the development of powerful LLM-based agents capable of automating the CFD workflow spectrum, from interpreting natural language specifications to configuring and executing simulations.
-This advancement represents a significant step toward making complex CFD simulations more accessible and efficient.
-OpenFOAM [18], in particular, stands out as the most popular open-source and free CFD software package, whose extensive capabilities, flexible architecture, and vibrant community support have made it a preferred choice for academic research, industrial applications, and consequently, a common framework for developing novel CFD automation agents.
-This accessibility has fostered innovation, leading to several pioneering initiatives exploring the integration of LLMs within the CFD domain.
-For example, MetaOpenFOAM [5] was introduced as an LLM-based multi-agent framework utilizing the MetaGPT [15] paradigm and RAG to automate CFD simulation tasks from natural language inputs, primarily demonstrated with OpenFOAM tutorial cases.
-Concurrently, OpenFOAMGPT [25] explored RAG-augmented LLM agents for OpenFOAM-centric CFD simulations, assessing model proficiency in tasks like zero-shot case setup and boundary condition adjustments.
-Foam-Agent [23] further advanced this area by proposing a multi-agent system with innovations such as hierarchical multi-index retrieval and dependency-aware file generation.
-Other research avenues have concentrated on fine-tuning LLMs with domain-specific datasets, like NL2FOAM [8], to enhance direct translation from natural language to executable CFD configurations, potentially using techniques like LoRA [16].
-Despite these significant strides in automating CFD pipeline components and reducing technical barriers, existing systems often struggle with the end-to-end management of complex simulations without human intervention.
+However, CFD implementation faces substantial barriers, requiring deep domain expertise [14] and often relying on expensive commercial software.
+Even experts invest significant time in tasks such as solver selection, model setup, mesh generation, boundary condition definition, and post-processing [15].
+These demands, combined with high computational costs, limit CFD accessibility for smaller organizations and hinder broader innovation [13].
+Thus, there is an urgent need for automated, intuitive, and affordable CFD solutions.
+To meet this need, we present ChatCFD, an AI-driven agent system that streamlines CFD workflows, facilitating automated scientific discovery in fluid mechanics and engineering by enabling rapid iteration and exploration of complex flow phenomena without extensive manual intervention.
+Recent progress in Artificial Intelligence (AI) has transformed the automation of sophisticated scientific processes.
+Large Language Models (LLMs), such as GPT [16], Gemini [17], and DeepSeek [18], alongside multi-agent frameworks like MetaGPT [19] and AutoGen [20], excel in natural language processing, code generation, and reasoning via chain-of-thought techniques [21].
+Frameworks like ReAct [22] synergize reasoning and acting in LLMs, while Reflexion [23] uses verbal reinforcement learning to improve agent performance through self-reflection.
+Toolformer [24] teaches LLMs to self-learn tool usage.
+Retrieval-Augmented Generation (RAG) [25] further bolsters these agents by incorporating domain-specific knowledge, reducing hallucinations, and improving domain adaptation for tasks like CFD automation.
+This synergy has enabled LLM-based agents to automate CFD pipelines, from interpreting user inputs to running simulations.
+OpenFOAM [26], a leading open-source CFD platform, supports this trend due to its flexibility and community backing.
+Notable efforts include MetaOpenFOAM [27], which automates simulations from natural language using RAG and MetaGPT, though limited to tutorial-level cases and struggling with complex geometries; OpenFOAMGPT [28], evaluating RAG-augmented LLMs for zero-shot setup and boundary adjustments, but facing difficulties in turbulence model modifications; Foam-Agent [29], introducing hierarchical retrieval for dependency-aware generation, yet constrained in handling unseen configurations; and NL2FOAM [30], fine-tuning LLMs for natural language-to-CFD translation using LoRA [31], but reliant on domain-specific datasets and less effective for diverse physical models.
+Despite these advances, existing systems often fall short in end-to-end automation for complex cases without human oversight, particularly due to a fundamental challenge: LLMs lack sufficient training on niche scientific tasks like OpenFOAM setup, necessitating expert-designed architectures to bridge this gap through structured reasoning, reflection, and knowledge integration.
 Existing CFD automation agents are often limited by their focus on rudimentary tasks, typically restricted to OpenFOAM tutorial-level examples, raising concerns about their generalization to complex, unseen cases.
-Three critical challenges remain unaddressed.
-First, describing CFD cases using brief natural language is often inadequate, as practical simulations require precise specifications beyond simplified verbal descriptions.
-If not given a detailed and professional task description, effective agents must interactively clarify user intent to define simulation requirements accurately.
-Second, handling intricate geometries poses significant difficulties, as mesh generation for real-world scenarios exceeds the capabilities of basic OpenFOAM utilities like “blockMesh” and demands specialized meshing tools.
-The inability to integrate user-defined or externally generated meshes limits these agents’ utility for diverse, practical CFD problems.
-Third, incorporating domain-specific knowledge into CFD agents remains challenging.
+This challenge primarily stems from the scarcity of domain-specific scientific coding corpora, such as OpenFOAM setups, requiring sophisticated multi-agent system designs to effectively incorporate physical knowledge into agent architectures, thereby enhancing LLMs’ capabilities in specialized scientific tool invocation.
+Recent representative work in this area includes SciToolAgent [32], which develops a knowledge graph-driven agent for integrating hundreds of scientific tools across biology, chemistry, and materials science, using graph-based RAG and safety-checking to automate workflows like protein engineering and chemical synthesis.
+While not directly focused on CFD, its approach to tool integration and domain adaptation offers valuable insights for fluid automation agents in handling coupled physics and complex simulations.
+Three critical challenges in CFD automation agents remain unaddressed.
+First, incorporating domain-specific knowledge into CFD agents remains challenging.
 Engineering-scale simulations require comprehensive initial and boundary condition specifications, necessitating domain-specific structural thinking and seamless integration of specialized domain expertise into the automation pipeline.
+Second, designing effective agent frameworks, especially for future multi-MCP agent collaborations, is essential to maximize complementary capabilities.
+Typical issues include whether we need to contain mesh generation in the agent scope, where real-world geometries exceed basic OpenFOAM tools like blockMesh and require specialized integration of external meshes.
+Third, enabling large-scale testing and evolution of fluid agents is crucial.
+Current reliance on simplistic natural language descriptions often leads to ambiguous case setups and hinders scalability.
+In contrast, CFD literature, comprising millions of papers with detailed descriptions and standard results, remains underutilized, impeding agent evolution.
 These shortcomings highlight a critical gap in current CFD automation frameworks.
@@
 The system comprises three core components: (1) an interactive chat interface for users to input case descriptions or upload mesh files, (2) a thinking system, the core decision-making module (detailed in Figure 2), and (3) a simulation engine that executes cases, collects error logs, and delivers final results.
-To address the limitations outlined above, this paper introduces ChatCFD, an automated agent system for configuring and executing CFD simulations within the OpenFOAM framework, as illustrated in Fig.
-1.
-ChatCFD enables users to define cases interactively via a chat interface or by uploading detailed technical files, such as well-documented research papers.
-Leveraging large language models (e.g., DeepSeek-R1, DeepSeek-V3), a sophisticated multi-agent architecture, and specialized OpenFOAM knowledge, ChatCFD autonomously manages the entire simulation workflow.
-It adeptly handles complex CFD setups through iterative trial-and-error, supporting comprehensive initial and boundary conditions, intricate physical models, and user-defined or externally generated mesh files.
-This architecture enables ChatCFD to address diverse, unseen scenarios and adapt various physical models, marking a significant advancement over prior agents.
-This paper presents the detailed architecture of the ChatCFD system and validates its modular effectiveness through experimental analysis.
-The paper is organized as follows: Section 2 details the system architecture, Section 3 provides experimental results demonstrating the effectiveness of its components, and Section 4 summarizes the key findings.
-An appendix includes illustrative examples and prompts to clarify ChatCFD’s operational functionality.
-2 Method
+To overcome these challenges, we introduce ChatCFD, an automated agent system for OpenFOAM simulations, as shown in Figure 1.
+ChatCFD processes multi-modal inputs (e.g., research papers, meshes) through an interactive interface, utilizing LLMs (DeepSeek-R1, DeepSeek-V3), multi-agent architecture, and OpenFOAM knowledge to manage the full workflow.
+It supports intricate setups via iterative refinement, handling diverse physical models and external meshes, surpassing prior agents in adaptability to unseen cases.
+This paper details ChatCFD’s architecture and validates its components experimentally.
+The structure is as follows: Section 2 describes the pipeline design, Section 3 presents results, and Section 4 summarizes findings.
+Appendices provide examples and prompts for operational clarity.
+2 ChatCFD Pipeline Design
 Figure 2: Architecture of the ChatCFD framework for automated CFD simulations, illustrating the four-stage workflow and agent structure.
-The stages are: (1) Knowledge Base Construction, creating a JSON database from OpenFOAM manuals and tutorials; (2) User Input Processing, enabling user interaction via natural language or document and mesh uploads; (3) Case File Initialization, generating OpenFOAM case files using the knowledge base; and (4) Simulation Execution and Error Reflection, running simulations, converting meshes with fluentMeshToFoam, and resolving errors (dimension mismatches, missing files, persistent errors, general issues) using RAG-based modules ReferenceRetriever and ContextRetriever.
+The stages are: (1) Knowledge Base Construction, creating a JSON database from OpenFOAM manuals and tutorials; (2) User Input Processing, enabling user interaction via natural language or document and mesh uploads; (3) Case File Generation, generating OpenFOAM case files using the knowledge base; and (4) Execution and Error Reflection, running simulations, converting meshes with fluentMeshToFoam, and resolving errors (dimension mismatches, missing files, persistent errors, general issues) using RAG-based modules ReferenceRetriever and ContextRetriever.
 The agent structure integrates DeepSeek-R1 and DeepSeek-V3 for intelligent processing, with iterative error correction.
@@
 A systematic analysis of OpenFOAM manuals and tutorials constructs a structured JSON database, cataloging essential parameters such as solver configurations, turbulence models, and file dependencies, providing a critical foundation for accurate case setup and error correction.
-Initially, preprocessed manual and tutorial data define the file dependency relationship, accessed during the RAG process to enhance the agent’s error correction ability.
-To ensure robust operation, strategic case filtering is implemented, removing OpenFOAM-specific headers, excluding cases with external dependencies or auxiliary folders to maintain independence, and filtering out cases with extensive non-uniform field definitions that reduce LLM response quality without contributing to configuration.
+Initially, preprocessed manual and tutorial data are used to define file dependency relationships, which are accessed during the RAG process to enhance the agent’s error correction capabilities.
+To ensure robust operation, strategic case filtering is implemented by removing OpenFOAM-specific headers and excluding cases with external dependencies or auxiliary folders.
+This maintains case independence, reducing the risk of degraded LLM response quality due to irrelevant configuration elements.
+For cases where non-uniform field definitions occupy significant file content and may exceed the dialogue context limits of large models, placeholder tokens such as __CELL_COUNT__,__FACE_COUNT__,__SCALAR_DATA_PLACEHOLDER__ are used to replace non-uniform field data.
+This approach enables the LLM to focus on critical settings, such as boundary and initial conditions, while ignoring non-uniform field configurations.
+As a result, it mitigates the reduction in the number of usable OpenFOAM tutorial cases caused by complex non-uniform field data.
 Filtering criteria and implementation details are provided in Appendix A.
@@
 For instance, the simpleFoam solver requires configuration files including system/fvSolution, system/controlDict, system/fvSchemes, 0/p, 0/U, constant/transportProperties, and constant/turbulenceProperties.
-Similarly, the kOmegaSST turbulence model requires system/fvSolution, system/controlDict, system/fvSchemes, 0/p, 0/U, 0/k, 0/omega, 0/nut, constant/turbulenceProperties, and constant/transportProperties.
+Similarly, the k-$\omega$ SST turbulence model requires system/fvSolution, system/controlDict, system/fvSchemes, 0/p, 0/U, 0/k, 0/omega, 0/nut, constant/turbulenceProperties, and constant/transportProperties.
 2.2 Stage 1: User Input Processing
@@
   • Case Selection: Through natural language interaction, the system guides users to select a target case, provides detailed specifications for verification, and confirms the selection.
-  • Mesh Integration: The system assists users in uploading mesh files, currently supporting Fluent-format .msh files.
+  • Mesh Integration: The system assists users in uploading mesh files, currently supporting Fluent format .msh files.
+For OpenFOAM default meshes, users must specify the path to the corresponding polyMesh directory.
 Upon completing this workflow, ChatCFD aggregates three essential components for downstream processing, as shown in Fig. 2:
@@
   • Mesh Data: Robust file transfer and path management ensure reliable delivery of mesh data to subsequent stages.
-2.3 Stage 2: Case File Initialization
-The CFD engineer layer, powered by the DeepSeek-R1 model, initializes OpenFOAM case files through a streamlined two-phase process, leveraging the preprocessed knowledge base.
-In the first phase, the system generates a configuration list of required case files by analyzing the case description from Stage 1 to determine the appropriate solver and turbulence model.
-It queries the knowledge base’s OpenFOAM dataset to compile a tailored list of necessary files (e.g., system/fvSolution, 0/p) for the solver-turbulence model pair, enhancing initialization accuracy and minimizing missing file errors.
-Such a procedure, based on an explicit manner, substantially enhances the quality of case file initialization and reduces the incidence of missing file errors.
-The second phase extracts detailed case setups from three sources: the configuration list, user prompt, and example library.
-The configuration list ensures precise case identification, while the prompt is segmented by section and filtered using the sentence-transformers/all-mpnet-base-v2 library with relevance thresholds.
-Hierarchical parameter extraction follows three steps:
-  • Boundary Condition Validation: The system identifies boundary types from the prompt, verifies compatibility with OpenFOAM v2406, and enforces strict naming conventions to avoid errors.
-For example, in the airfoil case shown in Fig.
-3, while the LLM may default to using a generic “inlet” type for the Inlet boundary based on common practice, the system ensures it correctly implements the specified “freeStream” boundary condition type to maintain fidelity with the original case setup.
-  • Field Value Specification: Initial and boundary condition values are extracted for all field files (e.g., 0/p, 0/nut, 0/nuTilda) listed in the configuration.
-  • Configuration File Generation: All required OpenFOAM files are generated, including field files (0/), physical properties (constant/, excluding polyMesh/), and control parameters (system/*).
-Step 3 leverages DeepSeek-R1’s reasoning capabilities to orchestrate the simultaneous generation of all required case configuration files while ensuring consistency across physical models and boundary conditions.
-For numerical configurations not explicitly specified in the research article, such as discretization schemes and solution algorithms defined in system/fvSchemes and system/fvSolution files, ChatCFD employs DeepSeek-R1’s advanced reasoning capabilities to construct appropriate configurations based on its understanding of CFD best practices and the physical requirements of the case.
-The agent maintains coherence between interdependent parameters while adapting numerical schemes to match the flow physics and solution requirements indicated by the article’s methodology.
-2.4 Stage 3: Simulation Execution and Error Reflection
-Stage 3 executes the CFD simulation and analyzes runtime logs for automated error correction through three sequential steps:
-  • Case Deployment: Case files from Stage 2 are deployed to the designated directory, and uploaded mesh files are converted to OpenFOAM format using fluentMeshToFoam, storing results in constant/polyMesh.
-  • Temporal Configuration: Temporal parameters in system/controlDict are set, running 10 iterations/time steps for steady-state or transient simulations, with outputs at steps 0, 5, and 10.
-Transient cases enforce a maximum CFL number of 0.6, with time steps of $1\times 10^{-5}$ s for incompressible flows and $1\times 10^{-8}$ s for compressible flows.
-  • Simulation Execution: The simulation runs using the specified OpenFOAM solver, with execution output logged to case_run.log for analysis.
-ChatCFD automates error detection and correction in OpenFOAM case files by analyzing runtime logs using RAG and two retrieval modules:
-The ReferenceRetriever module queries the preprocessed OpenFOAM knowledge database to select reference case files with matching or similar solvers and physical configurations.
-Typically, three reference files balance guidance and prompt complexity, ensuring contextually relevant corrections (e.g., distinguishing incompressible vs.
-compressible flows).
-The ContextRetriever module compiles all current case configurations into a structured JSON format, providing the DeepSeek-V3 model with comprehensive context to address errors like missing files, especially in physically coupled problems (Section 3.3).
-Upon detecting a runtime error, ChatCFD classifies it into four categories, detailed in Appendix B with examples:
-  • Dimension Mismatches: Inconsistent physical dimensions in field or configuration files are identified by DeepSeek-V3 through log pattern analysis.
-Corrections involve adjusting dimensions to match reference files from ReferenceRetriever, ensuring consistency.
-  • Missing Files: Missing case files, as reported in MetaOpenFOAM [5], are detected via log messages.
-DeepSeek-V3 identifies the missing file, ReferenceRetriever provides a template, and ContextRetriever supplies existing case details.
-DeepSeek-R1 generates the file, ensuring consistency in boundary patches, dimensions, and solver settings.
-  • Persistent Errors: Errors recurring across correction cycles are tracked by DeepSeek-R1.
-The problematic file is identified, and DeepSeek-V3 rewrites it using reference files from ReferenceRetriever and the original file content.
-  • General Errors: Errors like incorrect keywords or numerical schemes, comprising  50% of runtime issues, are addressed through a three-step process:
-    • Error Localization: DeepSeek-R1 analyzes the log and case files (via ContextRetriever) to pinpoint the erroneous file, constraining the search to relevant files to reduce LLM errors.
-    • Correction Proposal: DeepSeek-R1 generates step-by-step correction advice using the error log, erroneous file, and reference files from ReferenceRetriever, preserving initial and boundary conditions from Stage 1.
-    • Correction Application: DeepSeek-V3 applies the correction, adhering to the advice and maintaining dimensional consistency.
-2.5 LLM Configurations
-ChatCFD employs a dual-model approach, integrating DeepSeek-R1 and DeepSeek-V3 to handle the initiation and correction of case files in the workflow.
-The application of these two LLMs follows distinct principles.
-DeepSeek-R1 takes the lead in Stage 2, where it extracts comprehensive case setups from textual descriptions and configures the initial case files.
-Its capacity for in-depth text comprehension and error correction is also crucial in Stage 3, where it analyzes and rectifies general errors and generates any missing files.
-DeepSeek-V3’s contributions are concentrated in Stage 3, focusing on the correction of case files.
-It excels at managing highly patterned tasks, such as identifying dimensional and file-missing errors.
-Furthermore, DeepSeek-V3 handles specific corrections, such as adjusting field file dimensions, rebuilding case files to correct persistent errors, and amending files according to precise step-by-step instructions.
-This synergistic use of DeepSeek-R1 and DeepSeek-V3 significantly enhances ChatCFD’s overall performance, particularly in correcting case files.
-While DeepSeek-R1 excels in Stage 3 with its strong reasoning and planning for long-text comprehension and comprehensive file configuration, its tendency to hallucinate or generate inaccuracies became a hurdle during development.
-These hallucinations led to problems such as misidentifying dimensional errors and inserting unwanted markdown formatting into corrected files, issues that extensive prompt engineering alone could not resolve.
-To address these challenges, ChatCFD adopted a strategic hybrid method.
-In the case file correction module, DeepSeek-V3 now manages the majority of error corrections, which has markedly increased the success rate of these corrections.
-DeepSeek-R1 retains responsibility for resolving file-missing errors, as this task requires a holistic analysis of all configuration files for consistency—a strength of DeepSeek-R1’s content processing.
-A similar division applies to error type identification: DeepSeek-V3 handles pattern-based errors like dimension and file-missing issues, while DeepSeek-R1 tackles complex errors, such as repetitive and common problems, which necessitate broader content analysis.
-For DeepSeek-V3, the thought prompting by Wu et al.
-[35] in 2024 is used to equip DeepSeek-V3 with some thinking abilities.
-An auxiliary function then helps to extract the final answer from DeepSeek-V3’s response.
-The thought prompting text instructs the model to respond comprehensively and in detail, and to first write down its thought process after the phrase "Here is my thought process:" before providing its actual response after "Here is my response:".
-An example of thought prompting is given in Appendix C.
-Furthermore, ChatCFD maintains a log of all Question-Answer (QA) histories with the LLMs for each CFD case.
-These QA logs are valuable for analyzing the LLM’s behavior as it configures CFD cases within the OpenFOAM framework.
-The experimental validation detailed in Section 3 was conducted on the VolcEngine platform [31] from March 1 to March 12, 2025.
-The pricing structure for the models on this platform, used for evaluating token usage and cost, is as follows: for DeepSeek-R1, the cost is \$0.00055 (0.004 RMB) per 1,000 input tokens and \$0.0022 (0.016 RMB) per 1,000 output tokens; for DeepSeek-V3, the cost is \$0.00021 (0.0015 RMB) per 1,000 input tokens and \$0.00082 (0.006 RMB) per 1,000 output tokens.
-All token usage and cost calculations mentioned in Section 3 are based on this pricing structure from the VolcEngine platform.
+2.3 Stage 2: Case File Generation
+The CFD engineer layer, powered by the DeepSeek-R1 and DeepSeek-V3 models, initializes OpenFOAM case files through a streamlined three-phase process that leverages the preprocessed knowledge base.
+2.3.1 Phase 1: File Configuration Synthesis
+The system analyzes the case description from Stage 1 (User Input Processing) to identify suitable solvers and models, thereby generating a list of required case file configurations.
+This involves explicitly matching dependencies for solvers, models, and related settings to facilitate accurate file configuration.
+The workflow proceeds as follows:
+  • Relevant cases are retrieved based on the solver and model, yielding a corresponding file configuration list from which non-essential files (e.g., those for visualization or mesh generation) are removed.
+These refined cases are then input to the DeepSeek-V3 model to produce an optimized file configuration list.
+  • In the absence of direct matches, cases are aligned by solver type.
+For additional models, such as turbulence models, necessary files are inferred from analogous cases, synthesized into a new configuration list, verified for solver compatibility, and finalized via DeepSeek-V3.
+This approach integrates solver and physical model structures with case synthesis, enhancing initialization precision, minimizing file omissions, and substantially improving overall case file quality.initialization.
+2.3.2 Phase 2: Case Setup Extraction
+Detailed case configurations are derived from three primary sources: the file configuration list, user prompts, and the case library.
+The configuration list prevents omissions or redundancies in the case structure, while user prompts enable filtering of segmented case description chunks using the sentence-transformer/all-mpnet-base-v2 library and a similarity threshold.
+The DeepSeek-R1 model implements a hierarchical three-step extraction:
+  • Mesh Boundary Condition Extraction: Boundary names and types are extracted via regular expressions and tools like pyFoam, forming a persistent dictionary for subsequent physical field generation and error handling.
+  • Boundary Condition Validation: Boundary conditions are identified from filtered chunks, checked for compatibility with OpenFOAM-v2406, and subjected to rigorous naming and format validation.
+  • Physical Field File Value Setup: Field files (e.g., 0/p, 0/nut, 0/nutTilda) are identified from the configuration list.
+Using prior boundary data, OpenFOAM-compliant templates are created, encompassing internal and boundary field settings differentiated by boundary names and associated keywords.
+2.3.3 Phase 3: Configuration Validation and Refinement
+For unspecified parameters, such as discretization schemes and solver algorithms in system/fvSolution and system/fvSchemes, the DeepSeek-R1 model applies advanced reasoning grounded in CFD best practices and case-specific physics to formulate coherent configurations.
+This maintains parameter interdependencies and ensures alignment with the outlined flow dynamics and solver demands.
+A comprehensive validation and correction mechanism further refines the generated files.
+The system scrutinizes structures in the system and constant directories, confirming proper dependencies and completeness through cross-references with library exemplars.
+Particular emphasis is placed on physical field dimensions to mitigate LLM limitations in interpreting such constraints.
+For example, variables like p and alphat—which differ dimensionally between compressible and incompressible flows—are calibrated according to case type; ambiguous cases retain initial values.
+This process curtails errors like dimensional inconsistencies, bolsters initial file robustness, and diminishes iterative corrections in Stage 3.
+2.4 Stage 3: Error Correction and Reflection
+Stage 3 of the ChatCFD pipeline employs a modular architecture for automated error correction and iterative refinement in OpenFOAM case configurations.
+The following modules facilitate robust error handling:
+  • ReferenceRetriever: This module retrieves reference files from the preprocessed OpenFOAM knowledge base, matching solver and model specifications from benchmark tutorial cases.
+If no exact match is found, it prioritizes solver-compatible files.
+To balance guidance and prompt complexity, two reference files are selected to aid error detection and correction, leveraging the LLM’s few-shot learning to enhance ChatCFD’s accuracy in error localization and resolution.
+  • ContextRetriever: This module compiles current case configurations into a structured JSON format, providing detailed file content and modification trajectories over a specified period, tailored to other modules’ needs.
+It supports targeted reflection and correction, particularly for errors involving physical coupling (e.g., pressure-density interactions).
+  • Error Localization Module: Comprising two components: (1) DeepSeek-V3 rapidly identifies suspicious files based on error messages, reducing the analysis burden; (2) DeepSeek-R1 performs detailed reasoning to pinpoint erroneous files by comparing their content against the case library, leveraging its advanced inference capabilities for precise localization.
+  • Error Correction Module: This module addresses file interdependencies in OpenFOAM through a three-step process: (1) DeepSeek-V3 identifies files requiring coordinated modifications to resolve coupling-related errors; (2) DeepSeek-R1 proposes actionable corrections by integrating error details, related file content, simulation requirements, and benchmark tutorial cases; (3) DeepSeek-V3 applies these corrections, ensuring compliance with OpenFOAM’s file format standards.
+  • Reflection Module: Activated during persistent errors, this module collects error messages and file modification histories, formulating reflections in the format: ”For situation A, I considered B but overlooked C; next, I will apply D to resolve the issue.” These reflections are stored as reflection blocks within a <reflexion> tag in the reflection history, enhancing iterative error correction by integrating with localization and correction modules.
+Errors detected during simulation are categorized into two types:
+  • General Errors: These, comprising over 70% of runtime errors, include incorrect keywords, formatting issues, or floating-point errors.
+Correction involves a streamlined workflow: suspicious file detection (DeepSeek-V3), error confirmation (DeepSeek-R1), retrieval of related tutorial files (ReferenceRetriever), proposal of corrections, and file modification.
+For missing file errors, DeepSeek-V3 directly generates the required file using reference cases, followed by dimensional validation to ensure quality.
+  • Persistent Errors: ChatCFD leverages short-term memory (recent error messages and file modification histories) and long-term memory (reflection histories) to address recurring issues.
+Reflection histories, stored as structured insights, enhance the system’s ability to adapt and resolve complex errors iteratively.
+2.5 LLM Architecture and Configuration
+ChatCFD employs a dual-model architecture, integrating the DeepSeek-R1 and DeepSeek-V3 large language models (LLMs) to execute a comprehensive CFD workflow for case generation and error correction.
+Each model is strategically deployed based on its strengths to optimize performance across Stages 2 and 3.
+The DeepSeek-R1 model, with its superior text comprehension and reasoning capabilities, is utilized for complex, high-stakes tasks.
+In Stage 2, it extracts critical simulation parameters (e.g., boundary conditions, initial conditions) from case descriptions and generates complete initial case files.
+In Stage 3, it pinpoints error-causing files, proposes detailed correction strategies, and conducts reflection on persistent errors.
+However, DeepSeek-R1’s propensity for hallucinations—such as inserting unnecessary Markdown formatting, which can disrupt error correction—poses challenges.
+Despite mitigation efforts using prompt engineering and pydantic for structured outputs, these issues persist.
+Additionally, its inference time increases significantly with longer prompts, making it less efficient for routine tasks.
+The DeepSeek-V3 model excels in instruction following and rapid response, making it ideal for structured, low-complexity tasks.
+In Stage 2, it performs quick file structure validation, while in Stage 3, it identifies suspicious or missing files and generates corrected file content based on provided instructions.
+To address DeepSeek-V3’s limited reasoning capacity, tasks are simplified into structured formats (e.g., providing configuration file lists to narrow analysis scope), and the model is prompted to justify its actions (e.g., explaining why a file is suspicious).
+This enhances its effectiveness while maintaining computational efficiency.
+This dual-model approach leverages DeepSeek-R1’s reasoning for complex CFD tasks and DeepSeek-V3’s efficiency for structured operations, ensuring robust performance across diverse case types (e.g., benchmark tutorial cases, literature-derived cases) while minimizing computational costs.
+ChatCFD maintains a comprehensive log for each CFD case, capturing all question-answer interactions and actions taken during reflection-based error correction.
+These logs are invaluable for analyzing the behavior of the DeepSeek-R1 and DeepSeek-V3 models within the OpenFOAM framework, particularly in identifying limitations in handling complex CFD configurations.
+The recorded reflection iterations, which document corrective actions for persistent errors, provide critical insights into scenarios where ChatCFD’s performance is suboptimal, guiding targeted improvements in the pipeline.
+Additionally, a concise summary of reflections for recurring errors is maintained, offering a deeper understanding of ChatCFD’s behavior in complex scenarios.
+These logs have the potential to form a novel database, serving as a valuable resource for enhancing LLM-based agents in addressing high-complexity CFD problems within the MCP framework.
+The experimental validations described in Section 3 were conducted on the VolcEngine platform [33], which provides the pricing structure for token usage and computational costs.
+For DeepSeek-R1, input tokens cost \$0.00055 per 1,000 (0.004 RMB), and output tokens cost \$0.0022 per 1,000 (0.016 RMB).
+For DeepSeek-V3, input tokens cost \$0.00021 per 1,000 (0.0015 RMB), and output tokens cost \$0.00082 per 1,000 (0.006 RMB).
+All token consumption and cost metrics reported in Section 3 are based on this pricing, enabling precise evaluation of ChatCFD’s computational efficiency across benchmark tutorial cases, perturbed variant cases, and literature-derived cases.
 3 Results and discussion
-The performance of ChatCFD was assessed through validation experiments on basic CFD examples and two complex cases, using published literature as prompts.
-ChatCFD achieved error-free configurations in 30-40% of cases and operational success in 60-80%, marking a significant advance in automated CFD setup.
-The system demonstrated robust handling of complex configurations, particularly those with diverse turbulence models, highlighting its adaptability to varied physical scenarios.
-Analysis revealed two key performance factors: the extent of physical coupling in the flow and the availability of comprehensive knowledge about relevant physical models.
-These findings elucidate the capabilities and limitations of AI-driven CFD automation, contributing to the advancement of automated CFD workflows.
-3.1 CFD cases for Experiments
-Figure 3: Demonstration of the two tested CFD cases, including grids of (a) the Naca0012 case [29] and (b) the Nozzle case [36], and boundary types of (c) the Naca0012 case and (d) the Nozzle case.
-We selected two representative test cases for simulation: an incompressible flow and a compressible flow case (illustrated in Figure 3).
-These cases were chosen based on the comprehensive documentation available in their source articles, which provided detailed specifications of the solver configurations, turbulence models, and mesh characteristics essential for accurate CFD reproduction.
-For the incompressible flow case, we utilized the work of Sun et al.[29], who conducted simulations of a NACA0012 airfoil at 10° angle of attack employing simpleFoam solver with the SpalartAllmaras turbulence model.
-For the compressible flow case, we referenced Yu et al.[36], who performed simulations of a nozzle with pressure ratio 3.
-The rhoCentralFoam and SpalartAllmaras model are used in the Nozzle case.
-Given the documented challenges in modifying OpenFOAM turbulence models using LLM, as reported by Pandey et al.[25], we extended our investigation to include additional experiments with the NACA0012 case implementing various turbulence models.
-Table 1 presents a comprehensive overview of all case configurations and experimental iterations conducted in this study.
-High-quality computational meshes for these experiments were generated using commercial software.
-The experimental design featured a structured testing protocol: Cases 1 and 5 were subjected to an extensive model ablation analysis, with 10 iterations performed for each of the five system configurations.
-In contrast, Cases 2 through 4 were evaluated using only the complete system configuration, also with 10 iterations each.
-This systematic methodology yielded a total of 130 ChatCFD experimental runs, forming a robust dataset for subsequent analysis.
+The performance of ChatCFD was rigorously evaluated through a series of validation experiments encompassing three distinct categories of CFD cases: (i) 205 benchmark tutorial cases drawn from OpenFOAM tutorials and the OpenFOAM wiki, serving as standardized references for foundational validation; (ii) 110 perturbed variants, derived by systematically altering key parameters (e.g., boundary conditions, solver settings, or physical properties) in the benchmark cases to assess robustness and sensitivity; and (iii) 2 advanced literature-derived cases, directly prompted from published research articles to test real-world applicability in complex, unseen scenarios.
+ChatCFD demonstrated an operational success rate of 82.1% across the 315 benchmark and perturbed cases, with success defined as error-free configuration and execution leading to converged simulations.
+For the advanced literature-derived cases, success rates ranged from 60% to 80%, reflecting the challenges of handling intricate, domain-specific configurations without prior exposure.
+These results represent a substantial advancement in LLM-driven CFD automation, particularly in reducing the expertise barrier for setup and execution within the OpenFOAM framework.
+3.1 Performance Evaluation on OpenFOAM Basic Cases
+Figure 3: Comparison of success rates across three CFD agents (ChatCFD, MetaOpenFOAM, Foam-Agent) for 205 benchmark and 110 perturbed OpenFOAM tutorial cases.
+(a) Success rates by case category.
+(b) Distribution of test cases across categories.
+(c) Overall success rate comparison.
+To evaluate ChatCFD, a comprehensive set of test cases was curated from OpenFOAM tutorials and the OpenFOAM wiki.
+Cases involving complex mesh generation techniques were filtered out, resulting in 205 benchmark tutorial cases.
+From these, 11 cases were selected and systematically perturbed by modifying their descriptions and adjusting parameters (e.g., boundary conditions, solver settings, or physical properties), yielding an additional 110 perturbed variant cases.
+This approach ensures both foundational validation and robustness testing under varied conditions.
+Figure 3 illustrates the performance of ChatCFD compared to MetaOpenFOAM and Foam-Agent across these cases, showing success rates by category, case distribution, and overall performance.
+Success is defined as the generation of error-free case files leading to converged simulations.
+As shown in Figure 3(c), ChatCFD achieves an overall success rate of 82.1%, significantly outperforming MetaOpenFOAM (6.2%) and Foam-Agent (42.3%).
+Figure 3(a) highlights ChatCFD’s consistent advantage across case categories, demonstrating its superior performance and generalization capability.
+[…]
+Figure 4: Performance statistics for different agents across 205 benchmark tutorial cases.
+(a) Average token consumption per case.
+(b) Distribution of token consumption.
+(c) Average number of reflection iterations per case.
+(d) Average token consumption per reflection iteration.
+(e) Average computational cost (in monetary terms).
+(f) Distribution of reflection iteration ratios, excluding zero and limit-reaching cases.
+The performance of ChatCFD, MetaOpenFOAM, and Foam-Agent was evaluated through detailed metrics on 205 benchmark tutorial cases from OpenFOAM, focusing on token consumption, reflection iterations, and computational cost.
+Figure 4 presents a comparative analysis of these metrics, highlighting ChatCFD’s efficiency and robustness in automated CFD workflows.
+As shown in Figure 4(a), ChatCFD exhibits the lowest average token consumption per case, approximately half that of Foam-Agent, with MetaOpenFOAM ranking second.
+This efficiency underscores ChatCFD’s high success rate combined with reduced costs, making it practical for large-scale CFD applications.
+While ChatCFD and MetaOpenFOAM have comparable total token consumption, Figure 4(e) reveals a significant cost disparity, with ChatCFD’s computational cost being approximately 1.5 times lower than MetaOpenFOAM’s.
+This is primarily due to ChatCFD’s strategic use of the cost-efficient DeepSeek-V3 model (with a cost half that of DeepSeek-R1) for specific pipeline stages, as illustrated in Figure 4(b).
+In contrast, MetaOpenFOAM’s lower success rate necessitates frequent error corrections, generating additional file content and inflating costs.
+Figure 4(c) shows that MetaOpenFOAM requires the highest average number of reflection iterations per case, followed by Foam-Agent, with ChatCFD requiring the fewest.
+However, Figure 4(d) indicates that MetaOpenFOAM consumes fewer tokens per reflection iteration due to its simpler error-correction approach, which dilutes the high token cost of initial case setup across multiple iterations.
+Conversely, ChatCFD and Foam-Agent employ more sophisticated error-reflection strategies.
+Foam-Agent, for instance, maintains an “error correction trajectory” that logs modifications to file $i$ at iteration $j$, supplemented by contextual data from related files and reference cases.
+This increases its per-iteration token consumption.
+ChatCFD, while similarly providing contextual data, optimizes efficiency by maintaining both short-term modification trajectories and long-term reflection summaries.
+This dual approach enables precise filtering of irrelevant information and concise summarization of critical context, significantly reducing token consumption during LLM interactions compared to Foam-Agent, as evidenced in Figure 4(d).
+Figure 4(f) depicts the distribution of reflection iteration ratios, defined as the number of reflections divided by the maximum allowed iterations, reflecting each agent’s assessment of case complexity.
+The x-axis represents the reflection ratio, and the y-axis indicates the proportion of cases at each complexity level.
+Higher curves indicate stronger error-correction capabilities at a given reflection ratio.
+All three agents exhibit a rapid decline in solved cases followed by saturation, suggesting that increasing reflection iterations beyond a certain point yields diminishing returns.
+For Foam-Agent, with a reflection limit of 49 iterations, the curve flattens at approximately 50% (around 25 iterations), indicating that additional reflections beyond this threshold rarely resolve errors.
+ChatCFD’s curve consistently lies above the others, demonstrating more effective reflections that are likelier to correct errors.
+However, even for ChatCFD, reflections beyond 75% of the limit (approximately 22 iterations) often fail to resolve remaining errors, highlighting a current limitation in LLM-based CFD agents where complex case errors require advanced knowledge integration or alternative strategies beyond iterative reflection.
+3.2 Evaluation and Analysis on Literature-Derived CFD Cases
+Figure 5: Visualization of two literature-derived CFD cases.
+(a) Computational mesh for the NACA0012 airfoil case [34].
+(b) Computational mesh for the Nozzle case [35].
+(c) Boundary condition types for the NACA0012 case.
+(d) Boundary condition types for the Nozzle case.
+To evaluate ChatCFD’s ability to handle complex, real-world scenarios, two representative literature-derived cases were selected: an incompressible flow case and a compressible flow case, as illustrated in Figure 5.
+These cases were chosen for their comprehensive documentation in the source literature, providing detailed specifications for solver configurations, turbulence models, and mesh characteristics critical for accurate CFD reproduction.
+The incompressible flow case, based on Sun et al.
+[34], involves a NACA0012 airfoil at a 10° angle of attack, simulated using the simpleFoam solver with the Spalart-Allmaras turbulence model.
+The compressible flow case, drawn from Yu et al.
+[35], models a nozzle with a pressure ratio of 3, employing the rhoCentralFoam solver and the Spalart-Allmaras model.
+Given reported challenges in modifying OpenFOAM turbulence models using large language models (LLMs), as noted by Pandey et al.
+[28], additional experiments were conducted to explore robustness across various turbulence models (e.g., $\epsilon, k$-$\omega$ SST).
+Table 1 provides a comprehensive summary of all case configurations and experimental iterations.
+High-fidelity computational meshes were generated to ensure accuracy.
+The experimental protocol was structured as follows: Cases 1 (NACA0012) and 5 (nozzle) underwent an extensive model ablation study, with 10 iterations per case across five system configurations to assess sensitivity to model variations.
+Cases 2 through 4, using different turbulence models, were evaluated based on the full ChatCFD system configuration, with 10 iterations each, to focus on performance under optimal settings.
+This rigorous methodology resulted in 130 experimental runs, forming a robust dataset for analyzing ChatCFD’s performance in complex, literature-derived scenarios.
+The results, detailed in subsequent analyses, demonstrate ChatCFD’s capability to adapt to sophisticated CFD configurations, complementing its strong performance on benchmark tutorial and perturbed variant cases.
 Table 1: Summary of Test Cases for ChatCFD Validation
@@
 simpleFoam with
-SpalartAllmaras
+Spalart-Allmaras
 10 runs $\times$ 5 system
@@
 simpleFoam with
-kOmegaSST
+k-$\omega$ SST
 10 runs $\times$ the complete
@@
 simpleFoam with
-kEpsilon
+k-$\epsilon$
 10 runs $\times$ the complete
@@
 simpleFoam with
-RNGkEpsilon
+RNGk-$\epsilon$
 10 runs $\times$ the complete
@@
 hePsiThermo and
-SpalartAllmaras
+Spalart-Allmaras
 10 runs $\times$ 5 system
 configurations
-3.2 Model Ablation Analysis
-To systematically evaluate the contributions of different components and strategies, ablation experiments were performed using five distinct system configurations.
-These configurations vary in terms of comprehensive paper interpretation detailed in Stage 2 (Initiating Case Files), the sophistication of error handling logic, and the utilization of retrieval modules (ReferenceRetriever and ContextRetriever) in Stage 3 (Correcting Case Files):
-  • Configuration A (Baseline System): Utilizes a simplified parameter extraction process instead of the three-step process in Stage 2.
-In Stage 3, error correction is limited to a basic “general error” pathway, with both ReferenceRetriever (tutorial-based retrieval) and ContextRetriever (current case retrieval) modules deactivated.
-  • Configuration B (ReferenceRetriever with Basic Error Handling): Maintains the simplified parameter extraction (Stage 2) and basic ’general error’ handling (Stage 3) of Configuration A, but activates the ReferenceRetriever module to inform corrections with tutorial case data.
-  • Configuration C (Multi-Category Errors, ReferenceRetriever, Simplified Extraction): Employs the simplified parameter extraction in Stage 2.
-However, Stage 3 incorporates the full multi-category error correction logic (addressing dimensional, missing file, repetitive, and general errors) and activates the ReferenceRetriever module.
-The ContextRetriever module remains deactivated.
-  • Configuration D (Multi-Category Errors, ReferenceRetriever & ContextRetriever, Simplified Extraction): Also uses the simplified parameter extraction in Stage 2.
-Stage 3 employs the full multi-category error correction logic, supported by both ReferenceRetriever and ContextRetriever modules.
-  • Configuration E (Complete System): Represents the most comprehensive setup.
-It features comprehensive paper interpretation in Stage 2.
-In Stage 3, it utilizes the full multi-category error correction logic, augmented by both ReferenceRetriever and ContextRetriever modules.
-Figure 4: Success rates of ChatCFD in reproducing NACA0012 Case 1 and Nozzle Case 5 across five distinct Configurations A to E.
-The horizontal axis delineates these configurations, while the vertical axis quantifies the success rates.
-A 10-step operation indicates the CFD case successfully advanced 10 time steps or executed 10 iteration steps.
-Accuracy denotes precise case configuration in accordance with the source article’s specifications.
-Figure 4 illustrates the performance of ChatCFD across the five distinct system Configurations A through E for two benchmark test cases: NACA0012 Case 1 and Nozzle Case 5.
-Two primary metrics were employed for evaluation: the 10-step operational success rate, signifying the capability to complete ten simulation steps (time steps for transient cases or iterations for steady-state cases) without critical errors, and the accurate configuration success rate, indicating that the CFD case was set up in exact accordance with the specifications detailed in the source research article.
-The 10-step criterion was adopted to enable rapid assessment of operational stability for the system’s reflection mechanism, as target simulation durations are often unspecified in the literature.
-It was observed that configurations meeting this 10-step milestone were generally free of major operational faults, though minor discrepancies in boundary conditions parameters from the article’s description could still occur.
-For cases achieving accurate configuration, the simulation results demonstrated good agreement with published data, as detailed in Appendix D.
+3.2.1 Ablation Study on ChatCFD System Configurations
+To systematically dissect the contributions of ChatCFD’s core components, ablation experiments were conducted across five distinct system configurations.
+These configurations vary in the complexity of case file generation (Stage 2, Section 2.3), the error-handling logic, and the integration of retrieval modules (ReferenceRetriever and ContextRetriever) in Stage 3 (Correcting Case Files).
+The configurations are defined as follows:
+  • Configuration A (Baseline System): Employs a simplified parameter extraction process in Stage 2, bypassing the three-step hierarchical extraction.
+In Stage 3, error correction is limited to a basic ’general error’ pathway without ReferenceRetriever (tutorial-based retrieval).
+  • Configuration B (Baseline with ReferenceRetriever) and Simplified Setup Extraction: Retains the simplified Stage 2 extraction but enhances Stage 3 with the ReferenceRetriever module, enabling tutorial-based error correction within the general error pathway.
+  • Configuration C (Enhanced Setup and Reference Retrieval): Implements the full three-step setup extraction and configuration validation in Stage 2.
+Stage 3 includes general error correction augmented by the ReferenceRetriever module for improved accuracy.
+  • Configuration D (Full Error Reflection with Simplified Extraction): Uses the simplified Stage 2 extraction but incorporates both general and persistent error correction modules in Stage 3, supported by ReferenceRetriever and ContextRetriever for comprehensive error handling.
+  • Configuration E (Complete System): Represents the full ChatCFD pipeline, featuring comprehensive case setup extraction and validation in Stage 2 and complete error reflection modules in Stage 3, integrating both ReferenceRetriever and ContextRetriever for optimal performance.
+Figure 6: Success rates of ChatCFD across five system configurations (A–E) for literature-derived cases: NACA0012 (Case 1) and Nozzle (Case 5).
+The horizontal axis represents configurations, and the vertical axis quantifies success rates.
+The 10-step operational success rate indicates completion of 10 simulation steps (time steps for transient cases or iterations for steady-state cases) without critical errors.
+The accurate configuration success rate reflects precise adherence to source article specifications.
+Figure 6 illustrates ChatCFD’s performance across Configurations A through E for two literature-derived cases: NACA0012 (Case 1) and Nozzle (Case 5).
+Two metrics were evaluated: (1) the 10-step operational success rate, defined as the ability to complete 10 simulation steps (time steps for transient cases or iterations for steady-state cases) without critical errors, and (2) the accurate configuration success rate, indicating exact compliance with the solver, turbulence model, and boundary condition specifications in the source articles.
+The 10-step criterion enables rapid assessment of operational stability, particularly for the reflection mechanism, as literature often lacks specified simulation durations.
+Configurations achieving this milestone were generally free of major operational faults, though minor deviations in boundary condition parameters could persist.
+Cases meeting the accurate configuration criterion exhibited simulation results in strong agreement with published data, as detailed in the appendix.
+These results underscore ChatCFD’s robustness in handling complex, literature-derived configurations, complementing its performance on benchmark tutorial and perturbed variant cases.
 For the NACA0012 Case 1, the data in Figure 6 show a clear progression in the 10-step operational success rate with increasing system sophistication.
 The baseline Configuration A achieved a 20% success rate.
-This rate systematically improved to 40% for Configuration B, 60% for Configuration C, and reached 80% for Configuration D, which benefited from enhanced retrieval capabilities.
 This trend highlights the positive impact of incrementally integrating retrieval modules and OpenFOAM-specific knowledge.
-However, despite its high operational success, Configuration D failed to achieve any accurate configuration success, being unable to precisely replicate the case setup from the source article.
-This crucial gap was addressed by Configuration E, which incorporates comprehensive article interpretation and parameter extraction as part of its advanced Stage 2 processing.
+However, despite its high operational success, Configuration D failed to achieve any accurate configuration success, unable to precisely replicate the case setup from the source article.
+This crucial gap was addressed by Configuration E, which incorporates comprehensive article interpretation and setup extraction as part of its advanced Stage 2 processing.
 With Configuration E, the accurate configuration success rate for the NACA0012 case improved significantly to 40%, while maintaining the 80% operational success rate.
-The comprehensive article interpretation module in Stage 2 is integral to achieving accurate CFD configurations, particularly for cases involving non-standard boundary conditions.
-For instance, the NACA0012 case, as depicted in Figure 3, employs several specialized OpenFOAM boundary conditions (freestream, freestreamVelocity, and freestreamPressure) that deviate from common practices.
-Analysis of the LLM’s internal reasoning processes indicated that, without robust article interpretation capabilities and strict adherence to documented specifications, the model tends to default to more conventional boundary conditions.
-This can lead to erroneous configurations, such as using standard inlet-type conditions for inlet velocities and pressureOutlet-type conditions for outlet pressures, instead of the requisite specialized freestream conditions.
-This tendency for conventional settings also extends to numerical parameters, including inlet velocity values.
-To mitigate these issues, the comprehensive article interpretation module was developed in Stage 2, incorporating a three-step process for analyzing articles.
-This process emphasizes the precise extraction of case specifications and ensures adherence to OpenFOAM’s inherent boundary condition types, as detailed in Section 2.
-The introduction of this module yielded substantial improvements: while the 80% success rate for the 10-step operational benchmark was maintained, the system’s capability to accurately replicate documented case configurations increased dramatically from 0% to 40%.
-This significant enhancement underscores the efficacy of the refined interpretation strategy in correctly identifying and implementing specialized case requirements.
-In the Nozzle Case 5, as depicted in Figure 4, the introduction of the ContextRetriever module led to a substantial improvement in the 10-step operational success rate.
-Configurations lacking ContextRetriever (i.e., up to Configuration C) exhibited a 0% success rate for this metric.
-The integration of ContextRetriever in Configuration D elevated this rate to 60%.
-This marked improvement is attributed to ContextRetriever’s enhanced capability to diagnose and address complex coupling-related errors (refer to Appendix B for examples), a notable limitation in earlier configurations.
-For instance, when encountering thermophysical parameter errors in OpenFOAM (such as those related to $\Psi$, the compressibility), previous configurations often misidentified the error source as the constant/thermodynamicProperties file.
-However, such errors frequently originate from inconsistencies in the equation of state specifications within boundary condition entries in field files (e.g., 0/p or 0/T).
-The ContextRetriever module overcomes this by performing a comprehensive analysis across all relevant case directories (0/, constant/, and system/), thereby enabling more accurate error localization and effective correction strategies.
-The differential performance of the ContextRetriever module when applied to the NACA0012 Case 1 and the Nozzle Case 5 can be attributed to fundamental distinctions in their underlying flow physics.
-NACA0012 Case 1, an incompressible flow simulation employing the pisoFoam solver, exhibits a comparatively weak coupling between pressure and velocity fields.
-In contrast, Nozzle Case 5, a compressible flow scenario utilizing the rhoCentralFoam solver, is characterized by strong interdependencies among density, temperature, and pressure, governed by the equation of state.
-The advanced analytical capabilities of ContextRetriever, particularly in discerning inter-file dependencies and identifying errors related to coupled physics, proved notably more effective for the Nozzle case due to its inherent complexities.
-While still beneficial, the module’s impact was less pronounced for the NACA0012 case, which features a less coupled system.
-This observed variance underscores the proficiency of ContextRetriever in managing comprehensive physical information across a spectrum of problem complexities.
-Advancing from Configuration D to E for the Nozzle Case 5 preserved the 60% 10-step operational success rate while notably enhancing the accurate configuration success rate from 0% to 30%.
-This pattern of improvement in configuration precision, achieved by incorporating comprehensive article interpretation in Stage 2, mirrors the observations from the NACA0012 case, indicating a consistent benefit of the complete system configuration across different flow regimes.
-Figure 5: Results for NACA0012 Case 1 using Configurations A to E, showing averages of 10-step operational success runs.
-Metrics shown include: number of reflection iterations (#reflection), total token consumption for DeepSeek-R1 and DeepSeek-V3 models (R1/V3 total token), execution cost (Cost \$), and average token consumption per reflection round for both models (R1/V3 token/#reflection)
+Analysis of the DeepSeek-R1 model’s internal reasoning revealed that, without robust article interpretation and strict compliance with documented specifications in Stage 2, the LLM often defaults to generic boundary conditions and simplified parameters, such as standard inlet-type conditions for inflow velocities or pressureOutlet-type conditions for outflow pressures, instead of the specialized freestream conditions required for the NACA0012 airfoil.
+These erroneous configurations underscore the critical role of comprehensive case setup extraction and validation in Stage 2, which Configuration E leverages to ensure fidelity to complex CFD requirements, enhancing ChatCFD’s performance in complicated scenarios.
+For the Nozzle Case 5, Figure 6 demonstrates a significant enhancement in the 10-step operational success rate with the integration of the ContextRetriever module.
+Configurations A through C, lacking ContextRetriever, achieved a 0% success rate for this metric, defined as completing 10 simulation steps (time steps for transient cases or iterations for steady-state cases) without critical errors.
+Configuration D, incorporating ContextRetriever, improved this rate to 60%.
+This advancement stems from ContextRetriever’s ability to diagnose complex coupling-related errors within the MCP framework, addressing a key limitation of earlier configurations.
+For example, errors in thermophysical parameters, such as compressibility ($\Psi$), were often misattributed to the constant/thermophysicalProperties file in prior configurations.
+In reality, these errors frequently arise from inconsistencies in equation of state specifications within boundary condition entries in field files (e.g., 0/p, 0/T).
+By analyzing all relevant directories (0/, constant/, system/), ContextRetriever enables precise error localization and effective correction, enhancing performance for complex cases.
+The differential impact of ContextRetriever between the NACA0012 Case 1 and Nozzle Case 5 reflects their distinct flow physics.
+The NACA0012 case, an incompressible flow simulation using the pisoFoam solver, exhibits weak pressure-velocity coupling, reducing the need for advanced error handling.
+Conversely, Nozzle Case 5, a compressible flow scenario with the rhoCentralFoam solver, involves strong interdependencies among density, temperature, and pressure, governed by the equation of state.
+The ContextRetriever module’s advanced analysis of inter-file dependencies proves particularly effective for the Nozzle case’s coupled physics, yielding greater performance gains compared to the less coupled NACA0012 case.
+This highlights ContextRetriever’s proficiency in managing complex physical interactions across diverse CFD scenarios.
+Advancing to Configuration E, which integrates comprehensive article interpretation in Stage 2, maintained the 60% 10-step operational success rate for Nozzle Case 5 while increasing the accurate configuration success rate from 0% to 30%.
+This metric reflects precise adherence to source specifications.
+Consistent with findings for the NACA0012 case, Configuration E’s enhanced setup extraction and validation ensure greater configuration fidelity across flow regimes, reinforcing ChatCFD’s robustness in literature-derived cases.
+Figure 7: ChatCFD’s performance metrics for accurate configuration of literature-derived NACA0012 Case 1 and Nozzle Case 5.
+Metrics include: number of reflection iterations (#reflection), total token consumption for DeepSeek-R1 and DeepSeek-V3 models (R1/V3 total token), execution cost (Cost \$), and average token consumption per reflection round for both models (R1/V3 token/#reflection).
+The performance disparity between the literature-derived NACA0012 Case 1 and Nozzle Case 5 arises from two key factors: (1) the increased complexity of managing numerous files with intricate content in compressible flows, and (2) the stringent requirement for consistency across coupled thermophysical models (e.g., 0/p, 0/T, constant/thermophysicalProperties).
+The Nozzle case, a compressible flow simulation using rhoCentralFoam, involves strong coupling among pressure, density, temperature, and the equation of state, necessitating precise dimensional and physical alignment across governing equations and material properties.
+This contrasts with the NACA0012 case, an incompressible flow simulation using pisoFoam with weaker pressure-velocity coupling, which imposes fewer constraints.
+Additionally, an implicit bias in LLM training data, predominantly focused on incompressible flow scenarios, leads to erroneous defaults in compressible cases.
+For instance, in OpenFOAM, incompressible flows use kinematic pressure ($p_{k}$, dimensions $L^{2}T^{-2}$)—dynamic pressure divided by density—while compressible flows require absolute thermodynamic pressure ($p$, dimensions $ML^{-1}T^{-2}$).
+The DeepSeek-R1 model’s tendency to apply incompressible conventions to compressible cases introduces dimensional errors, such as incorrect pressure settings, requiring additional reflection iterations for correction.
+Figure 7 quantifies these challenges: the Nozzle case requires over twice the reflection iterations, triple the DeepSeek-R1 tokens, and six times the DeepSeek-V3 tokens compared to the NACA0012 case, resulting in approximately triple the execution cost.
+The disproportionate DeepSeek-V3 token increase (6x vs.
+3x for DeepSeek-R1) reflects frequent dimensional inconsistencies and persistent errors, necessitating more invocations of the costlier DeepSeek-V3 model.
+Figure 7(c) confirms this, showing a DeepSeek-V3 token multiplier of  2.7x per reflection for the Nozzle case, compared to  2x for DeepSeek-R1, relative to the NACA0012 case.
+These metrics highlight the computational burden of complex flow simulations and underscore ChatCFD’s ability through advanced error correction.
+3.2.2 Impact of Turbulence Models on ChatCFD Performance
+Figure 8 presents ChatCFD’s performance across four turbulence models—Spalart-Allmaras (SA), $\epsilon, k$-$\omega$ SST, and RNG k-$\epsilon$—applied to the literature-derived NACA0012 cases (Cases 1 to 4).
+The analysis focuses on simulations achieving accurate configuration per published specifications, excluding those with only basic operational success (e.g., 10-step operational success rate).
+All results were obtained using the complete ChatCFD system (Configuration E).
+As shown in Figure 8(a), success rates for accurate configuration vary significantly: SA at 40%, k-$\omega$ SST at 30%, k-$\epsilon$ at 20%, and RNG k-$\epsilon$ at 10%.
+This demonstrates ChatCFD’s ability to handle diverse turbulence models, addressing LLM challenges noted in OpenFOAMGPT [28] through structured knowledge integration and RAG.
+Performance metrics in Figure 8(a,b) reveal consistent profiles for SA, k-$\epsilon$, and k-$\omega$ SST, with an average of  6 reflection iterations, comparable DeepSeek-R1 token usage, and costs.
+In contrast, RNG k-$\epsilon$ requires  23.5 reflections—nearly 4x higher—elevating token consumption and costs.
+Figure 8(c) attributes these disparities to configuration file token counts and model prevalence in OpenFOAM tutorials.
+While file sizes differ modestly ( 1,800 tokens for SA vs.
+2,000 for others), tutorial distribution is more influential: SA, k-$\epsilon$, and k-$\omega$ SST appear in 28+ tutorials each, versus one for RNG k-$\epsilon$.
+RNG k-$\epsilon$’s performance deficit stems from: (1) its rarity in engineering applications, leading to sparse LLM training data and poor zero-shot configuration; and (2) limited ReferenceRetriever efficacy due to few OpenFOAM tutorials, hindering guidance for parameters in constant/turbulenceProperties, system/fvSchemes, and system/fvOption.
+This necessitates greater reliance on the LLM’s limited intrinsic knowledge, reducing performance relative to common models.
+Figure 8: ChatCFD’s average performance for accurate configuration of NACA0012 Cases 1 to 4 across turbulence models (Spalart-Allmaras, $\epsilon, k$-$\omega$ SST, RNG k-$\epsilon$).
+Metrics shown include: accuracy (success rate of precise configuration according to literature), number of reflection iterations (#reflection), total token consumption of the DeepSeek-R1 model (R1 total token), execution cost (Cost \$), token count of configuration files (#token of config files), and the number of cases incorporating each turbulence model in OpenFOAM tutorials (#Case of model in OF tutorial).
+3.3 Future Potential of ChatCFD in Multi-Agent CFD Workflows
+Building on the insights from ChatCFD’s performance in benchmark tutorial, perturbed variant, and literature-derived cases, ongoing work explores its potential as a foundational component in advanced engineering workflows.
+Beyond functioning as a standalone tool for CFD automation, ChatCFD is being adapted through the Model Context Protocol (MCP), a standardized framework for integrating specialized AI agents into a cohesive multi-agent system.
+MCP facilitates seamless communication and data exchange among agents by defining structured protocols for context sharing, enabling ChatCFD to collaborate with other intelligent tools to address complex, interdisciplinary engineering tasks.
+As illustrated in Figure 9, this multi-agent workflow enables an end-to-end design cycle for CFD applications.
+The process begins with a 3D modeling agent that generates a three-dimensional geometry from a natural language description or a simple image, such as a sketch or photograph.
+A subsequent meshing agent automatically discretizes the geometry into a high-fidelity computational mesh, leveraging tools like commercial mesh generators or open-source alternatives (e.g., blockMesh, snappyHexMesh).
+This mesh is then passed to the ChatCFD agent, which autonomously configures solver parameters, boundary conditions, and turbulence models (e.g., simpleFoam with Spalart-Allmaras or rhoCentralFoam with k-$\omega$ SST) to execute the CFD simulation, computing key performance metrics such as drag coefficients or pressure distributions.
+Finally, an optimization agent analyzes the simulation results, providing insights for iterative design refinement, such as adjusting geometry or boundary conditions to minimize drag or enhance flow stability.
+This MCP agent is available to use and can be found online (https://www.bohrium.com/apps/designagent0812).
+This integrated approach highlights ChatCFD’s transformative potential beyond simulation automation.
+By serving as a standardized computational module within a collaborative multi-agent ecosystem, ChatCFD can contribute to broader engineering design processes, from conceptual design to performance optimization.
+The MCP framework ensures interoperability, allowing ChatCFD to adapt to diverse tasks, such as aerodynamics, heat transfer, or multiphase flow simulations, by interfacing with agents specialized in geometry generation, mesh refinement, or post-processing.
+This vision positions ChatCFD as a cornerstone for future AI-driven engineering workflows, paving the way for scalable, interdisciplinary applications in scientific discovery and industrial innovation.
+Figure 9: ChatCFD integrated within a multi-agent workflow via the Model Context Protocol (MCP), enabling transformation from a single sentence or image to a complete CFD simulation.
+4 Conclusion
+This paper introduces ChatCFD, an innovative LLM-based CFD agent system for automating OpenFOAM simulations.
+By incorporating a multi-agent architecture, domain-specific knowledge bases, and advanced error reflection mechanisms, ChatCFD achieves end-to-end automation for reproducing literature results, including both incompressible (NACA0012 airfoil) and compressible (supersonic nozzle) flows.
+The simulations accurately reproduce velocity contours, pressure coefficients, and schlieren images, aligning closely with experimental data.
+ChatCFD’s innovations include: (1) an interactive multi-modal input mechanism integrating PDFs, meshes, and dialogue with physical constraints, providing a ChatGPT-like interface for guided literature analysis, case selection, and parameter specification, reducing operational barriers; (2) specialized AI agents leveraging CFD expertise for task interpretation, feature library construction via OpenFOAM preprocessing, and iterative reflection-refinement using DeepSeek-R1 for reasoning and DeepSeek-V3 for precise modifications; and (3) a robust error reflection system with ReferenceRetriever (tutorial templates) and ContextRetriever (case configurations) for real-time error classification and resolution.
+Experimental validations on 315 basic cases (205 benchmark tutorials, 110 perturbed variants) yield an 82.1% operational success rate, outperforming MetaOpenFOAM (6.2%) and Foam-Agent (42.3%), with lower success in multiphase/combustion due to interdisciplinary complexity.
+On literature-derived cases, ChatCFD achieves 60–80% success, with ablation studies showing comprehensive interpretation boosts accurate configuration from 0% to 40% for NACA0012 and 30% for Nozzle, while ContextRetriever improves operational success from 0% to 60% in Nozzle by resolving coupled-physics errors.
+Turbulence model analyses reveal 40% success for common models (e.g., Spalart-Allmaras) vs.
+10% for rare ones (e.g., RNG k-$\epsilon$), attributed to data scarcity.
+Physics coupling studies indicate compressible flows demand tripled resources due to interdependencies, and LLM bias toward incompressible setups causes dimensional errors in complex cases.
+These findings highlight ChatCFD’s robustness, efficiency, and cost-effectiveness, establishing benchmarks for AI-driven CFD.
+ChatCFD effectively addresses the critical gaps in CFD automation agents.
+First, it integrates domain-specific knowledge through structured RAG and expert-designed agents, mitigating LLM training scarcity to enable precise boundary and initial condition specifications.
+Second, its modular, MCP-compatible design supports collaborative multi-agent networks, integrating specialized meshing agents to handle complex geometries beyond blockMesh.
+Third, by leveraging extensive literature-derived corpora with detailed specifications and benchmarks, it will facilitate scalable testing and iterative agent evolution, harnessing millions of CFD papers for refinement and validation in the future.
+These advancements accelerate scientific discovery in fluid mechanics and engineering by automating hypothesis testing, parameter optimization, and multiphysics simulations.
+ACKNOWLEDGEMENTS
+This work is sponsored by the National Natural Science Foundation of China, Grant No. 92470127, and the Overseas Postdoctoral Talents Program in Guangdong Province.
+AUTHOR DECLARATIONS
+Conflict of Interest
+The authors declare no conflicts of interest.
+Author Contributions
+E Fan: Methodology, Software, Data curation, Formal analysis, Resources, Writing - original draft.
+Kang Hu: Methodology, Software, Data curation, Formal analysis, Resources.
+Zhuowen Wu: Methodology.
+Jiangyang Ge: Data curation, Formal analysis.
+Jiawei Miao: Software, Resources.
+Yuzhi Zhang: Conceptualization, Software, Resources.
+He Sun: Resources, Writing - Review.
+Weizong Wang: Resources, Writing - Review, Funding acquisition.
+Tianhan Zhang: Conceptualization, Methodology, Formal analysis, Resources, Writing - Review & Editing, Visualization, Supervision, Funding acquisition
+Data Availability
+The data supporting the findings are available at https://github.com/ConMoo/ChatCFD.
+Appendix A Appendix: detailed analysis of two literature-derived cases
 Figure 10 presents a detailed analysis of the NACA0012 Case 1 performance across the five distinct Configurations (A through E), based on averaged results from 10-step operational success runs.
@@
 The subsequent stabilization of DeepSeek-V3 token usage from Configuration C to E reflects the maturation of the reflection module’s design, achieving optimal operational efficiency without necessitating additional DeepSeek-V3 module invocations.
+Figure 10: Results for NACA0012 Case 1 using Configurations A to E, showing averages of 10-step operational success runs.
+Metrics shown include: number of reflection iterations (#reflection), total token consumption for DeepSeek-R1 and DeepSeek-V3 models (R1/V3 total token), execution cost (Cost \$), and average token consumption per reflection round for both models (R1/V3 token/#reflection)
 Figure 11: Results for Nozzle Case 5 using Configurations D and E, showing averages of 10-step operational success runs.
@@
 This consistency underscores the efficacy of the system’s architectural design, indicating that the comprehensive paper interpretation module detailed interpretation component operates as a distinct layer that augments overall system performance without adversely affecting the fundamental reflection mechanisms.
-3.3 Performance degradation in configuring physical-coupled CFD cases
 Table 2: Statistical averages of accurate configured cases for Naca0012 Case 1 and Nozzle Case 5.
 Metrics Naca0012, Case 1 Nozzle, Case 5
-Physical models Flow type: Incompressible Solver: simpleFoam Turbulence: SpalartAllmaras Flow type: Compressible Solver: rhoCentralFoam Turbulence: SpalartAllmaras Thermo model: hePsiThermo
+Physical models Flow type: Incompressible Solver: simpleFoam Turbulence: Spalart-Allmaras Flow type: Compressible Solver: rhoCentralFoam Turbulence: Spalart-Allmaras Thermo model: hePsiThermo
 Number of configuration files 9 12
@@
 DeepSeek-V3 call count 17.4 35.2
-Figure 7: ChatCFD’s average performance per execution when accurately configuring NACA0012 Case 1 and Nozzle Case 5.
-Metrics shown include: number of reflection iterations (#reflection), total token consumption for DeepSeek-R1 and DeepSeek-V3 models (R1/V3 total token), execution cost (Cost \$), and average token consumption per reflection round for both models (R1/V3 token/#reflection).
-The model ablation analysis in Section 3.2 highlights significant performance variations among configurations A through E when applied to two distinct test cases: the incompressible NACA0012 Case 1 and the compressible Nozzle Case 5.
-These discrepancies are attributed to fundamental differences in their underlying flow physics, with subsequent discussion detailing the influence of physical coupling—a key characteristic of compressible flow prominent in the Nozzle case—on ChatCFD’s performance.
-The Nozzle case utilizes the rhoCentralFoam solver, which inherently manages strong interdependencies between pressure, temperature, and density through its thermophysical models.
-Its compressible flow solution also dictates supersonic or subsonic conditions based on local velocity and sound speed, which are coupled with thermodynamic properties (e.g., pressure, temperature).
-In contrast, the NACA0012 case employs the simpleFoam solver, featuring minimal such coupling.
-As shown in Table 2, these differing physical complexities translate to substantially different configuration requirements.
-The Nozzle case demands more configuration files with greater token content and three additional thermophysical-related files (0/T, 0/alphat, constant/thermodynamicProperties).
-Furthermore, achieving a precise configuration for this compressible flow scenario necessitates over twice the DeepSeek-R1 and DeepSeek-V3 model calls compared to the incompressible NACA0012 case.
-Section 3.2 results demonstrate that the strong physical coupling in the Nozzle case correlates with diminished success rates.
-Figure 4 illustrates this, showing that the highly coupled Nozzle case consistently achieves lower 10-step success and precise configuration rates across ten experimental runs.
-This performance degradation is attributed to two primary factors: the increased complexity of managing a larger number of files with more extensive content, and the stringent requirement for maintaining consistency across interconnected thermophysical models and related parameters (e.g., in files such as 0/p, 0/T, and constant/thermodynamicProperties).
-The greater physical complexity of the Nozzle case translates directly to substantially higher LLM consumptions than NACA0012 Case 1 (Figure 7).
-Specifically, the Nozzle simulation requires over double the reflection iterations for convergence, consumes approximately triple the total DeepSeek-R1 tokens, and uses six times the DeepSeek-V3 tokens.
-These escalated resource needs lead to the Nozzle case incurring roughly triple the operational costs.
-Such pronounced differences in resource utilization directly quantify the greater computational burden imposed by the strong physical coupling in compressible flow simulations relative to their incompressible counterparts.
-Notably, the DeepSeek-V3 token consumption for the Nozzle case increased by a factor of approximately six, substantially exceeding the DeepSeek-R1 token increase of approximately threefold.
-Analysis of case execution QA histories reveals this disparity stems from a higher incidence of dimensional inconsistencies and persistent errors during Nozzle case configuration, which in turn necessitated more frequent invocations of the DeepSeek-V3 model.
-Figure 7(c) corroborates this, showing that the per-reflection DeepSeek-V3 token multiplier for the Nozzle case (about 2.7x) is obviously larger than its DeepSeek-R1 multiplier (about 2x) compared to the NACA0012 case.
-The analysis of dimensional errors reveals a notably higher frequency in compressible flow scenarios, such as the Nozzle case, attributable to two fundamental factors.
-First, compressible flow physics imposes inherently more stringent requirements for achieving consistent setups; this involves not only the dimensional consistency and intricate coupling of primary fields like pressure, density, and velocity, but critically, their coherent integration with a thermodynamic model (such as an equation of state).
-This model, which interrelates these variables, often via temperature and specific energy, demands precise dimensional and physical alignment across all governing equations and material property definitions, a level of rigor that significantly elevates the complexity compared to typical incompressible simulations.
-Second, and significantly, the predominant focus on incompressible flow applications in typical CFD scenarios appears to have created an implicit bias in LLM training data.
-This bias manifests as a tendency for the LLM to default to incompressible flow configurations, even when handling compressible cases.
-A pertinent example of how this bias can degrade compressible flow configurations is the differing treatment of pressure dimensions: in OpenFOAM, incompressible flows often implement kinematic pressure $p_{k}$, which is dynamic pressure over density and has dimensions of $L^{2}T^{-2}$ whereas compressible flows use absolute dynamic (thermodynamic) pressure (p) with dimensions $ML^{-1}T^{-2}$.
-An LLM incorrectly applying the incompressible pressure dimension convention to a compressible case due to this training bias would introduce a fundamental dimensional error.
-Consequently, when LLMs encounter compressible flow scenarios, these embedded assumptions can lead to dimensional erros, exemplified by the pressure dimension mismatch, that require additional correction cycles to resolve.
-3.4 Effects of turbulence models on ChatCFD’s performance
-Figure 8: ChatCFD’s performance when applying four turbulence models to the NACA0012 Cases 1 to 4, where the results represent the average outcomes of cases precisely configured according to the paper specifications.
-The turbulence models include SpalartAllmaras, kEpsilon, kOmegaSST, and RNGkEpsilon.
-Metrics shown include: accuracy (success rate of precise configuration according to literature), number of reflection iterations (#reflection), total token consumption of the DeepSeek-R1 model (R1 total token), execution cost (Cost \$), token count of configuration files (#token of config files), and the number of cases incorporating each turbulence model in OpenFOAM tutorials (#Case of model in OF tutorial).
-Figure 8 presents a detailed performance analysis of ChatCFD across four distinct turbulence models—Spalart-Allmaras (SA), kEpsilon, kOmegaSST, and RNGkEpsilon—applied to the NACA0012 benchmark (Cases 1 to 4).
-The evaluation focuses on simulations that achieved accurate configuration in accordance with published literature specifications, thereby excluding instances that only demonstrated rudimentary operational functionality (i.e., 10-step operational success).
-All experimental results reported herein were obtained using Configuration E, representing the complete agentic system.
-As illustrated in Figure 8(a), the analysis reveals distinct success rates for accurate case configuration across the four turbulence models.
-The SA model achieved the highest success rate at 40%, followed by the kOmegaSST model (30%) and the kEpsilon model (20%).
-The RNGkEpsilon model exhibited markedly lower performance, with a success rate of 10%.
-Notably, while OpenFOAMGPT [25] has underscored the difficulties LLMs encounter when modifying turbulence models during case configuration, ChatCFD’s capability to successfully implement all four models attests to its efficacy in addressing these challenges, thereby validating the proposed architectural design and implementation methodology.
-A detailed comparison of performance metrics, presented in Figure 8(a,b), delineates a clear behavioral divergence.
-The SA, kEpsilon, and kOmegaSST models displayed consistent performance profiles, characterized by an average reflection count of approximately 6, alongside comparable DeepSeek-R1 token utilization and associated computational expenditures.
-Conversely, the RNGkEpsilon model demanded substantially greater computational resources, evidenced by an average reflection count of 23.5—nearly fourfold higher than the other models—leading to correspondingly elevated token consumption and operational costs.
-To elucidate the origins of these performance disparities, an analysis of underlying factors was conducted, with results presented in Figure 8(c).
-This investigation considered both the token counts of configuration files and the prevalence of each turbulence model within the OpenFOAM tutorial suite.
-Although configuration file sizes exhibited only modest variations—the SA model requiring around 1,800 tokens compared to approximately 2,000 for the others—this minor difference is insufficient to account for the substantial performance gap.
-A more significant factor emerged from the distribution of tutorial cases: whereas the SA, kEpsilon, and kOmegaSST models were each featured in 28 or more tutorial examples, the RNGkEpsilon model was represented in merely one instance.
-The pronounced performance deficit observed for the RNGkEpsilon model is attributable to two principal factors.
-Firstly, its relative infrequency in practical engineering applications translates to a scarcity of representative examples in the LLM’s training dataset.
-This limitation impairs the LLM’s proficiency in configuring RNGkEpsilon cases effectively in a zero-shot manner (i.e., without RAG).
-Secondly, the model’s minimal presence within the OpenFOAM tutorials severely curtails the efficacy of the ReferenceRetriever module in sourcing appropriate reference configurations from this dataset.
-This is particularly detrimental for RNGkEpsilon-specific parameters, such as those found in the constant/turbulenceProperties file and associated numerical schemes within system/fvSchemes and system/fvOption.
-This dearth of tutorial-based guidance necessitates a greater dependence on the LLM’s intrinsic, yet less developed, understanding of the RNGkEpsilon model, culminating in diminished performance relative to more ubiquitously employed turbulence models.
-4 Conclusion
-This paper introduced ChatCFD, an innovative LLM-based CFD agent system designed to automate CFD simulations within the OpenFOAM framework.
-By integrating a multi-agent architecture, domain-specific knowledge bases, and advanced error correction mechanisms, ChatCFD successfully demonstrates the feasibility of end-to-end automation for both incompressible (NACA0012 airfoil) and compressible (supersonic nozzle) flow cases.
-Compared with antecedent CFD agent systems, such as MetaOpenFOAM [5] and OpenFOAMGPT [25], ChatCFD distinguishes itself through several key innovations:
-Firstly, ChatCFD advances the application paradigm of LLM agents in engineering simulation through an interactive guidance mechanism that innovatively integrates multimodal inputs—PDF files, mesh files, and interactive dialogue—with underlying physical constraints.
-This mechanism, featuring a dedicated mesh input module and a reinforced context-memory dialogue system, offers a ChatGPT-like interface.
-This interface systematically guides users through essential stages, including literature analysis, case selection, the uploading of user-defined meshes, and the interactive specification of CFD model parameters, thereby addressing the complexities inherent in practical CFD applications.
-Consequently, this approach substantially reduces operational hurdles compared to traditional command-line driven methodologies.
-Secondly, the system employs a collaborative framework of specialized AI agents that leverage deep CFD domain expertise.
-This strategy involves: (a) a comprehensive article interpretation module, utilizing an advanced reasoning LLM (e.g., DeepSeek-R1), for extracting CFD case parameters (e.g., boundary names, types, values) from academic papers, augmented by extensive OpenFOAM-specific boundary condition information to mitigate model hallucination, which is vital for precise CFD case replication; (b) preprocessing data from OpenFOAM tutorial cases to construct comprehensive feature libraries, thereby enriching the system’s understanding through RAG; and (c) a "global planning-local optimization" hierarchical framework.
-This framework strategically synergizes the advanced reasoning and text interpretation capabilities of one LLM with the high-fidelity, low-hallucination file modification strengths of another.
-For instance, when correcting configuration errors, DeepSeek-R1 first analyzes error messages and relevant files to pinpoint the erroneous file and suggest corrections, which DeepSeek-V3 then implements, ensuring both accuracy and reliability.
-Thirdly, ChatCFD features a powerful multi-category error correction mechanism with two specialized RAG modules: ReferenceRetriever (accessing tutorial templates) and ContextRetriever (collecting current case configurations).
-This feature actively monitors simulation outputs in real-time, dynamically selecting appropriate and effective treatments by classifying typical case configuration problems.
-These include, but are not limited to, dimensional inconsistencies, missing file errors, and recurring errors indicative of deeper setup issues, thereby systematically addressing critical logical consistency challenges frequently encountered by LLMs in complex CFD computations.
-Systematic ablation studies and case validations underscored the critical contributions of ChatCFD’s components.
-Comprehensive paper interpretation proved paramount for achieving high configurational accuracy, particularly for cases with rare or non-standard boundary conditions (e.g., “freestream” for NACA0012), improving accurate configuration success from 0% to 40% for the NACA0012 case and from 0% to 30% for the Nozzle case.
-The ContextRetriever module, which provides contextual awareness of current case files, was especially vital for diagnosing and rectifying complex, physically-coupled errors in compressible flow simulations; this boosted the Nozzle case’s 10-step operational success rate from 0% to 60%.
-Overall, the system achieved accurate configuration for approximately 30-40% of tested scenarios directly from literature, with higher operational success rates (e.g., 80% for NACA0012 with full system capabilities), establishing a valuable performance benchmark.
-Furthermore, the study quantified the cost-effectiveness of different system configurations.
-For instance, a 30% cost increase yielded a 40% improvement in accurate configuration for the NACA0012 case, justifying the investment for enhanced precision.
-These results highlight that while LLMs can automate aspects of CFD, their efficacy is significantly amplified when augmented with structured domain knowledge and robust error-handling mechanisms.
-Key findings indicate that ChatCFD’s performance is significantly impacted by the complexity and representation of CFD cases in its knowledge base.
-Configuring physically-coupled compressible flow scenarios, like the Nozzle case, drastically degrades performance compared to incompressible ones due to more numerous and intricate configuration files and the critical need for consistency across interdependent thermophysical parameters.
-This results in lower success rates for accurate configurations, substantially increased LLM calls, higher token consumption (especially for error correction by DeepSeek-V3), more reflection iterations, and tripled operational costs, with dimensional errors exacerbated by the inherent rigor of compressible physics and an LLM training bias towards incompressible flow conventions.
-Similarly, performance varies markedly with the chosen turbulence model; those well-represented in OpenFOAM tutorials (e.g., Spalart-Allmaras with 40% success) yield better results and lower resource use.
-In contrast, models like RNGkEpsilon, scarce in both general LLM training data and specific tutorial datasets (only one example), exhibit poor success (10%) and demand significantly more computational resources, underscoring how limited data representation hampers ChatCFD’s configuration proficiency and RAG effectiveness.
-ChatCFD pioneers computational automation, extending LLM applications to engineering challenges.
-By reducing expertise and software barriers, it democratizes access to advanced methodologies, fostering innovation.
-Its approaches—structured knowledge integration and context-aware RAG—offer insights for AI-driven engineering domains, such as simulations and optimizations.
-Future work includes expanding the knowledge base and developing standardized metrics to enhance automation, paving the way for AI-assisted scientific discovery.
-ACKNOWLEDGEMENTS
-This work is sponsored by the National Natural Science Foundation of China Grant No.
-92470127, 92270203, the Open Project of the National Key Laboratory of Scramjet Technology No.
-WDZC6142703202403, and the Overseas Postdoctoral Talents Program in Guangdong.
-AUTHOR DECLARATIONS
-Conflict of Interest
-The authors declare no conflicts of interest.
-Author Contributions
-E Fan: Methodology, Software, Data curation, Formal analysis, Resources, Writing - original draft.
-Weizong Wang: Resources, Funding acquisition.
-Tianhan Zhang: Conceptualization, Methodology, Formal analysis, Resources, Writing - Review & Editing, Visualization, Supervision, Funding acquisition
-Data Availability
-The data that support the findings of this study are available from the corresponding author upon reasonable request.
+Appendix B Appendix: CFD results of Cases 1 to 5
+Figure 12: Velocity magnitude contours for NACA0012 Cases 1 to 4 at 10° angle of attack using four turbulence models.
+(a) Spalart-Allmaras model, (b) k-$\epsilon$ model, (c) k-$\omega$ SST model, (d) RNGk-$\epsilon$ model.
+Figure 13: Comparison of pressure coefficients between CFD simulation and experimental data for NACA0012 airfoil at 10° angle of attack, using (a) Spalart-Allmaras model, (b) k-$\epsilon$ model, (c) k-$\omega$ SST model, and (d) RNGk-$\epsilon$ model.
+Symbols represent experimental results reported by Gregory et al.
+[37]
+Figures 12 and 13 show the velocity magnitude contours and surface pressure coefficient distributions for the NACA0012 airfoil Cases 1 to 4 at 10° angle of attack using different turbulence models.
+As can be seen, the velocity contours from the four cases are similar, and the pressure coefficient distributions agree well with the experimental results [37].
+Figure 14: (a) Experimental schlieren image [35] and (b) numerical schlieren image of Case 5 for the Nozzle case at nozzle pressure ratio = 3.
+Figure 14 illustrates the experimental and numerical schlieren images for the nozzle case at a nozzle pressure ratio of 3.
+The experimental results are reported in Ref.
+[35].
+The numerical results exhibit a strong correspondence with the experimental observations, accurately capturing key flow features.
+The figure clearly depicts the formation of a Mach stem, resulting from the intersection and reflection of shock waves from the ramp and flap.
+This interaction also reveals the characteristic Mach reflection, distinguished by its $\lambda$-shock wave structure.

### v3

@@
-ChatCFD: An LLM-Driven Agent for End-to-End CFD Automation with Domain-Specific Structured Reasoning
+ChatCFD: An LLM-Driven Agent for End-to-End CFD Automation with Structured Knowledge and Reasoning
 Abstract
 Computational Fluid Dynamics (CFD) is essential for advancing scientific and engineering fields, but is hindered by operational complexity, high expertise requirements, and limited accessibility.
-This paper introduces ChatCFD, an automated agent system for OpenFOAM simulations that processes multi-modal inputs (e.g., research papers, meshes) via an interactive interface, leveraging DeepSeek-R1 and DeepSeek-V3 large language models (LLMs), a multi-agent architecture, and OpenFOAM knowledge.
-Its four-stage pipeline, Knowledge Base Construction, User Input Processing, Case File Generation, and Execution and Error Reflection, enables iterative trial-reflection-refinement for intricate setups, supporting diverse physical models and external meshes.
-Validation on 205 benchmark tutorial cases, 110 perturbed variants, and 2 literature-derived cases demonstrates ChatCFD’s 82.1% operational success rate on basic cases, significantly outperforming MetaOpenFOAM (6.2%) and Foam-Agent (42.3%), and 60–80% on literature-derived complex cases, with turbulence model studies showing a 40% success rate for common models vs.
-10% for rare ones like RNGk-$\epsilon$.
-Physics coupling analyses reveal higher resource demands for multi-physics-coupled cases, while LLM bias toward simpler setups introduces persistent errors, such as dimensional inconsistency.
-Ablation studies highlight the efficacy of RAG-based modules and reflection mechanisms.
-By automating hypothesis testing and parameter exploration, ChatCFD accelerates scientific discovery in fluid mechanics and engineering, addressing LLM limitations through structured design and demonstrating strong potential as a modular component in MCP-based agent networks for collaborative multi-agent systems, paving the way for scalable AI-driven CFD innovation.
+This paper introduces ChatCFD, a LLM-driven agent system for end-to-end CFD automation.
+Powered by DeepSeek-R1/V3, a multi-agent architecture, structured OpenFOAM knowledge bases, precise error locator, and iterative reflection, ChatCFD dramatically outperforms prior systems.
+On 315 benchmark cases it attains 82.1% execution success (vs.
+6.2% MetaOpenFOAM, 42.3% Foam-Agent) and, crucially, 68.12% physical fidelity—the first rigorous metric capturing whether a runnable simulation is scientifically meaningful.
+A dedicated Physics Interpreter achieves 97.4% summary fidelity, exposing the striking LLM gap between fluent narration and enforcing dozens of tightly coupled physical constraints in executable code.
+Resource-efficiency analysis highlights ChatCFD’s practical advantage, achieving an average of 192.1k tokens and \$0.208 per case—half the tokens of Foam-Agent and 1.5× cheaper than MetaOpenFOAM.
+Ablation studies confirm structured domain knowledge and reasoning is indispensable: removing the Solver Template DB collapses accuracy to 48%, while the Error Locator Module proves the single most critical component.
+Flexibility experiments show autonomous solver selection across compressible/incompressible and steady/transient regimes (95.23% success) and turbulence-model switching (100% success), even on unseen configurations.
+By faithfully reproducing complex literature cases NACA0012 airfoil and supersonic nozzle with 60–80% end-to-end success where baselines fail entirely, ChatCFD establishes new benchmarks for AI-driven CFD automation.
+Physics coupling analyses reveal higher resource demands for multi-physics-coupled cases, while LLM bias toward simpler setups introduces persistent errors.
+ChatCFD’s modular, MCP-compatible design directly enables collaborative multi-agent networks and paves the way for scalable AI-driven CFD innovation.
 The code for ChatCFD is available at: https://github.com/ConMoo/ChatCFD.
@@
 1 Introduction
-Computational Fluid Dynamics (CFD) is a cornerstone technology in diverse scientific and engineering disciplines, including aerospace [1, 2], energy systems [3, 4], urban environment [5, 6, 7], combustion [8, 9], and biomedical applications [10, 11].
+Computational Fluid Dynamics (CFD) is a cornerstone technology in diverse scientific and engineering disciplines, including aerospace [28, 10], energy systems [15, 2, 36], urban environment [20, 17, 26], combustion [24, 39, 33], astrophysics[40], and biomedical applications [22, 9].
 It provides indispensable tools for simulating intricate fluid behaviors, thereby enabling design innovation and scientific discovery [3, 4].
@@
 2, the ChatCFD framework implements a comprehensive workflow to streamline simulation setup, execution, and analysis.
+Figure 3: Based on the knowledge base constructed from the contents such as the OpenFOAM manual and OpenFOAM tutorials, and how it functions within the ChatCFD framework
 2.1 Stage 0: Knowledge Base Construction
@@
 This study focuses on preprocessing OpenFOAM manuals and publicly accessible examples.
-A systematic analysis of OpenFOAM manuals and tutorials constructs a structured JSON database, cataloging essential parameters such as solver configurations, turbulence models, and file dependencies, providing a critical foundation for accurate case setup and error correction.
+Through a systematic analysis of OpenFOAM manuals and tutorials, we construct a structured JSON database that records key parameters such as solver configurations, turbulence models, and file dependencies, laying the foundation for accurate case setup and error correction.
+Additionally, a vector database is established to further enhance the understanding of OpenFOAM and to generate more precise OpenFOAM case descriptions.
+As shown in Fig.
+3, the database primarily comprises two components: a Structured Database in JSON format and a Vector Database for efficient retrieval.
+The Structured Database includes the following parts:
+  • File Dependency and Structure DB (db1): Designed to construct a hierarchical file structure model.
+This model is based on OpenFOAM tutorial cases and categorized primarily by solver type and turbulence model to store the necessary file structure for specific combinations.
+It also records the actual file structure of each tutorial case, which is utilized to construct and validate the space of permissible configurations.
+  • Boundary Condition DB (db2): Comprehensively catalogs all available boundary condition types within OpenFOAM.
+For each boundary condition, it precisely documents the required additional setting keywords and their configuration specifications, ensuring the completeness and validity of the boundary setup.
+  • Parameter Dimension DB(db3): Systematically records the standard dimensional configuration rules for all physical quantities in OpenFOAM.
+This database further defines the dimensional requirements and consistency constraints for relevant field variables under varying physical problems or specific solving environments.
+  • Solver Template DB (db4): Systematically stores the core file content and structured metadata of fundamental OpenFOAM tutorial cases.
+The metadata includes: the case’s designated solver, turbulence model, and physical problem category.
+This information serves as the key basis for achieving high-accuracy classification and rapid screening of tutorial cases, thus providing precise reference templates for automated case initialization and configuration.
+All configuration files are carefully parsed, converted to JSON format, and tagged with metadata (e.g., solver specifications and flow characteristics) for easy retrieval and application.
+OpenFOAM-specific header files are removed, and cases with external dependencies or auxiliary folders are excluded to maintain case autonomy and reduce the influence of irrelevant configuration elements on the LLM’s response quality.
+Furthermore, for cases where non-uniform field definitions occupy a large volume of file content (potentially exceeding dialogue context limits), placeholders (e.g., __CELL_COUNT__, __FACE_COUNT__, __SCALAR_DATA_PLACEHOLDER__) are used to replace the non-uniform field data.
+This allows the LLM to focus on critical settings (such as boundary and initial conditions) while mitigating the issue of a reduced case count due to complex data.
 Initially, preprocessed manual and tutorial data are used to define file dependency relationships, which are accessed during the RAG process to enhance the agent’s error correction capabilities.
@@
 If no exact match is found, it prioritizes solver-compatible files.
-To balance guidance and prompt complexity, two reference files are selected to aid error detection and correction, leveraging the LLM’s few-shot learning to enhance ChatCFD’s accuracy in error localization and resolution.
+To balance guidance and prompt complexity, two reference files are selected to aid error detection and correction, leveraging the LLM’s few-shot learning to enhance ChatCFD’s accuracy in error locator and resolution.
   • ContextRetriever: This module compiles current case configurations into a structured JSON format, providing detailed file content and modification trajectories over a specified period, tailored to other modules’ needs.
 It supports targeted reflection and correction, particularly for errors involving physical coupling (e.g., pressure-density interactions).
-  • Error Localization Module: Comprising two components: (1) DeepSeek-V3 rapidly identifies suspicious files based on error messages, reducing the analysis burden; (2) DeepSeek-R1 performs detailed reasoning to pinpoint erroneous files by comparing their content against the case library, leveraging its advanced inference capabilities for precise localization.
+  • Error Locator Module: Comprising two components: (1) DeepSeek-V3 rapidly identifies suspicious files based on error messages, reducing the analysis burden; (2) DeepSeek-R1 performs detailed reasoning to pinpoint erroneous files by comparing their content against the case library, leveraging its advanced inference capabilities for precise error location.
   • Error Correction Module: This module addresses file interdependencies in OpenFOAM through a three-step process: (1) DeepSeek-V3 identifies files requiring coordinated modifications to resolve coupling-related errors; (2) DeepSeek-R1 proposes actionable corrections by integrating error details, related file content, simulation requirements, and benchmark tutorial cases; (3) DeepSeek-V3 applies these corrections, ensuring compliance with OpenFOAM’s file format standards.
-  • Reflection Module: Activated during persistent errors, this module collects error messages and file modification histories, formulating reflections in the format: ”For situation A, I considered B but overlooked C; next, I will apply D to resolve the issue.” These reflections are stored as reflection blocks within a <reflexion> tag in the reflection history, enhancing iterative error correction by integrating with localization and correction modules.
+  • Reflection Module: Activated during persistent errors, this module collects error messages and file modification histories, formulating reflections in the format: ”For situation A, I considered B but overlooked C; next, I will apply D to resolve the issue.” These reflections are stored as reflection blocks within a <reflexion> tag in the reflection history, enhancing iterative error correction by integrating with error locator and correction modules.
 Errors detected during simulation are categorized into two types:
@@
 Reflection histories, stored as structured insights, enhance the system’s ability to adapt and resolve complex errors iteratively.
+2.5 Stage 4: Post-Processing
+To enhance the transparency and user intelligibility of the computation process, ChatCFD integrates a dedicated Physics Interpreter in the fourth stage of the workflow.
+This component, implemented using the DeepSeek-R1 model, automatically produces a technical summary immediately after the case is successfully generated and deemed executable.
+The output is presented in Markdown format, supplemented by structured expressions such as tables and file tree diagrams.
+Its content systematically covers multiple dimensions, including fundamental information, physical problem background, special configuration settings, file structure, and key file contents, thereby comprehensively ensuring the transparency of the simulated physical process for the user.
 2.6 LLM Architecture and Configuration
@@
 These results represent a substantial advancement in LLM-driven CFD automation, particularly in reducing the expertise barrier for setup and execution within the OpenFOAM framework.
-3.1 Performance Evaluation on OpenFOAM Basic Cases
-Figure 3: Comparison of success rates across three CFD agents (ChatCFD, MetaOpenFOAM, Foam-Agent) for 205 benchmark and 110 perturbed OpenFOAM tutorial cases.
-(a) Success rates by case category.
-(b) Distribution of test cases across categories.
-(c) Overall success rate comparison.
+3.1 Benchmarking Execution Success Rates Across Agents
 To evaluate ChatCFD, a comprehensive set of test cases was curated from OpenFOAM tutorials and the OpenFOAM wiki.
@@
 Figure 4(a) highlights ChatCFD’s consistent advantage across case categories, demonstrating its superior performance and generalization capability.
+Figure 4: Comparison of success rates across three CFD agents (ChatCFD, MetaOpenFOAM, Foam-Agent) for 205 benchmark and 110 perturbed OpenFOAM tutorial cases.
+(a) Success rates by case category.
+(b) Distribution of test cases across categories.
+(c) Overall success rate comparison.
 […]
@@
 However, even for ChatCFD, reflections beyond 75% of the limit (approximately 22 iterations) often fail to resolve remaining errors, highlighting a current limitation in LLM-based CFD agents where complex case errors require advanced knowledge integration or alternative strategies beyond iterative reflection.
-3.2 Evaluation and Analysis on Literature-Derived CFD Cases
+3.2 Physical Fidelity Evaluation and Semantic Error Analysis
+In Section 3.1, we assessed ChatCFD, MetaOpenFOAM, and Foam-Agent using success rate as the primary metric.
+While this allows rapid quantitative evaluation, successful execution does not ensure physical meaningfulness in CFD.
+To address this, we introduce physical fidelity as a core metric, evaluating whether executed simulations accurately capture user-specified physics.
+We also re-examine the link between run success and physical fidelity via agent-generated cases.
+We define physical fidelity through a three-tier protocol. A case qualifies only if it meets all criteria without simplifications that compromise the scientific objective or user requirements:
+  • Conditions Verification: Initial and boundary conditions match the user-described scenario.
+  • Model Confirmation: Physical properties, models, and parameters are appropriate and consistent.
+  • Flow Feature Comparison: Post-processed visualizations align with ground-truth flow characteristics.
+Using this protocol, we analyzed ChatCFD and Foam-Agent results (Fig.
+6).
+ChatCFD’s physical fidelity rate among runnable cases (phy|run) is lower than Foam-Agent’s, but its overall rate is higher, confirming superior performance.
+Note that phy|run, limited to runnable cases, may inflate for agents handling only simple setups.
+Figure 6: Physical fidelity evaluation of ChatCFD.
+(a) Success rates of ChatCFD and Foam-Agent under three criteria: run — execution success (simulation completes without crashing); $\midrun — phy$sical fidelity conditioned on execution success (proportion of runnable cases that are physically correct); phy — overall physical fidelity (proportion of all cases that are both runnable and physically meaningful).
+(b) Root-cause analysis of semantic failures in ChatCFD-generated cases that execute successfully but lack physical fidelity, categorized by error type (boundary/initial conditions, physical properties/models, numerical schemes, etc.).
+Additionally, we summarized the causes of physical incorrectness in runnable cases.
+The results, shown in Fig 6, indicate that incorrect settings of boundary conditions, initial conditions, and initial fields in multiphase flows are the main contributors, accounting for 60 % of the issues, as these require deep physical insight and geometric awareness.
+Such misconfigurations rarely trigger convergence failures, allowing apparent success that misleads agents.
+Similarly, omissions in setFieldsDict often evade detection, exacerbating infidelity.
+To further enhance transparency, we integrated a Physics Interpreter, evaluated using DeepSeek-V3 and GPT-4o for summary rationality.
+It achieved 97.4% fidelity, confirming accurate natural-language conveyance of the agent’s intended physics.
+The contrast between the near-perfect fidelity of the Physical Interpreter (97.4%) and the overall Physical Fidelity of only 68.12% is particularly interesting.
+This suggests that while a seemingly simple user request (e.g., "2-D flow around a cylinder at Re = 100") can be almost perfectly restated in natural language by the LLM, its faithful implementation in executable OpenFOAM code requires dozens of tightly interdependent, domain-specific settings (such as inlet velocity profile, turbulence suppression, kinematic viscosity, blockage ratio, spanwise boundary conditions, etc.).
+This highlights a key limitation of current LLMs: they excel at high-level linguistic narration but consistently struggle to enforce the complete chain of physical constraints within the generated executable code.
+3.3 Ablation Study on ChatCFD Architectural Components
+To quantitatively assess the contribution of the proposed database system and functional modules to the overall system performance, we conducted an ablation study using $20\%$ of the cases sampled from the OpenFOAM benchmark test set.
+It should be noted that the degree of influence exerted by different components will vary due to the inherent differences in task types.
+Consequently, for the specific task of literature replication, we will perform additional ablation experiments in Section 3.5.1 for further, dedicated analysis.
+The first four experiments were designed to evaluate the impact of the knowledge databases on ChatCFD’s performance.
+This knowledge base primarily influences Stage2 and Case File Generation and Error Correction and Reflection in Stage3.
+Simultaneously, the experimental groups "w/o mod1" (removal of the Error Locator Module) and "w/o mod2" (removal of the Reflection Module) were used to verify and quantify the efficacy of the error locator and reflection mechanisms, with their influence predominantly focused on Error Correction and Reflection in Stage3.
+The impact of the first four knowledge base ablation experiments on ChatCFD’s performance, ranked from highest to lowest influence, is as follows: Solver Template DB (db4, 48 %), Boundary Condition DB (db2, 57 %), Parameter Dimension DB (db3, 68 %), and File Dependency and Structure DB (db1, 73 %).
+The results show that the Solver Template DB is utilized across all operational stages of ChatCFD and is crucial for leveraging the Large Language Model’s few-shot learning capability to generate initially correct case files.
+Figure 7: Ablation study quantifying the contribution of individual prior-knowledge components and core modules to ChatCFD’s configuration accuracy.
+- w/o File Dependency DB (DB1): removes the structured JSON database that enforces mandatory file hierarchy and dictionary dependencies.
+- w/o Boundary Condition DB (DB2): disables the boundary-condition template database (no validated keyword requirements).
+- w/o Parameter Dimension DB (DB3): disables automatic dimension enforcement and file-format checks.
+- w/o Solver Template DB (DB4): removes all reference tutorial snippets and metadata.
+- w/o Error Locator Module (Module 1): disables the error-positioning module.
+- w/o Reflection (Module 2): disables the iterative reflection loop.
+- w/ all: full ChatCFD system (all four prior-knowledge databases + all function modules).
+Fig 7 shows that the removal of the Error Locator Module (w/o mod1) resulted in the lowest success rate.
+Here, we bypassed it and prompted DeepSeek-R1 to modify files based solely on error messages.
+The experimental results reveal two findings: 1) Accurate identifying of the erroneous file is a critical step for enhancing the agent’s efficiency during the reflection and correction phase (Stage 3); and 2) Relying only on error messages is ineffective.
+Its shortcomings are particularly pronounced when dealing with errors that do not explicitly point to the source file, such as dimensional errors.
+Removing the Reflection Module (w/o mod2) causes the smallest performance drop because it primarily benefits high-difficulty cases, which constitute only a minor fraction of the benchmark.
+In contrast, the other components—knowledge bases and Error Locator—impact cases across all difficulty levels.
+3.4 Flexibility Evaluation of Alternative OpenFOAM Configurations
+To evaluate ChatCFD’s ability to adapt to different physical assumptions and numerical settings, we designed two targeted experiments:
+  • Solver selection across regimes — Given only a textual description, ChatCFD autonomously determines the flow regime (compressible/incompressible, steady/transient) and selects the appropriate solver.
+  • Turbulence-model switching — With mesh and boundary conditions fixed, only the turbulence closure is changed, requiring automatic rewriting of constant/turbulenceProperties and related dictionaries.
+Fig 8 summarizes the results of Experiment 1.
+ChatCFD correctly identified the required solver in 20 of 21 cases, including the unseen “simpleCar” compressible-transient configuration (not present in the database).
+The sole failure occurred on “buoyantCavity” under compressible-transient assumptions, which currently exceeds ChatCFD’s supported scope.
+In the turbulence-switching experiment (Figure 9), ChatCFD successfully applied $\epsilon, k$-$\omega$ SST, and laminar closures to every case—including the originally laminar damBreak—demonstrating robust handling of non-default configurations.
+These results confirm ChatCFD’s strong generalization: it is not limited to tutorial-default settings but can reason about physically valid alternatives, significantly enhancing its practical utility for practical CFD exploration.
+Figure 8: Solver flexibility under varying physical regimes.
+Blue circles indicate the reference (tutorial) setting; green checks mark successful execution with the identified solver; red crosses denote failure.
+ChatCFD correctly selects the appropriate solver in 20/21 of cases, including unseen compressible-transient scenarios.
+Figure 9: Turbulence-model flexibility.
+ChatCFD successfully switches between $\epsilon, k$-$\omega$ SST, and other closures for all tested cases, including the originally laminar damBreak (100% success).
+Symbols follow Figure 8.
+3.5 Evaluation of CFD Cases Reproduced from Literature
 Figure 10: Visualization of two literature-derived CFD cases.
@@
 configurations
-3.2.1 Ablation Study on ChatCFD System Configurations
+3.5.1 Ablation Study on ChatCFD Configurations for Literature Reproduction
 To systematically dissect the contributions of ChatCFD’s core components, ablation experiments were conducted across five distinct system configurations.
@@
 In reality, these errors frequently arise from inconsistencies in equation of state specifications within boundary condition entries in field files (e.g., 0/p, 0/T).
-By analyzing all relevant directories (0/, constant/, system/), ContextRetriever enables precise error localization and effective correction, enhancing performance for complex cases.
+By analyzing all relevant directories (0/, constant/, system/), ContextRetriever enables precise error locator and effective correction, enhancing performance for complex cases.
 The differential impact of ContextRetriever between the NACA0012 Case 1 and Nozzle Case 5 reflects their distinct flow physics.
@@
 These metrics highlight the computational burden of complex flow simulations and underscore ChatCFD’s ability through advanced error correction.
-3.2.2 Impact of Turbulence Models on ChatCFD Performance
+3.5.2 Influence of Turbulence Models on ChatCFD Efficacy
 Figure 13 presents ChatCFD’s performance across four turbulence models—Spalart-Allmaras (SA), $\epsilon, k$-$\omega$ SST, and RNG k-$\epsilon$—applied to the literature-derived NACA0012 cases (Cases 1 to 4).
@@
 This paper introduces ChatCFD, an innovative LLM-based CFD agent system for automating OpenFOAM simulations.
-By incorporating a multi-agent architecture, domain-specific knowledge bases, and advanced error reflection mechanisms, ChatCFD achieves end-to-end automation for reproducing literature results, including both incompressible (NACA0012 airfoil) and compressible (supersonic nozzle) flows.
-The simulations accurately reproduce velocity contours, pressure coefficients, and schlieren images, aligning closely with experimental data.
-ChatCFD’s innovations include: (1) an interactive multi-modal input mechanism integrating PDFs, meshes, and dialogue with physical constraints, providing a ChatGPT-like interface for guided literature analysis, case selection, and parameter specification, reducing operational barriers; (2) specialized AI agents leveraging CFD expertise for task interpretation, feature library construction via OpenFOAM preprocessing, and iterative reflection-refinement using DeepSeek-R1 for reasoning and DeepSeek-V3 for precise modifications; and (3) a robust error reflection system with ReferenceRetriever (tutorial templates) and ContextRetriever (case configurations) for real-time error classification and resolution.
-Experimental validations on 315 basic cases (205 benchmark tutorials, 110 perturbed variants) yield an 82.1% operational success rate, outperforming MetaOpenFOAM (6.2%) and Foam-Agent (42.3%), with lower success in multiphase/combustion due to interdisciplinary complexity.
-On literature-derived cases, ChatCFD achieves 60–80% success, with ablation studies showing comprehensive interpretation boosts accurate configuration from 0% to 40% for NACA0012 and 30% for Nozzle, while ContextRetriever improves operational success from 0% to 60% in Nozzle by resolving coupled-physics errors.
-Turbulence model analyses reveal 40% success for common models (e.g., Spalart-Allmaras) vs.
-10% for rare ones (e.g., RNG k-$\epsilon$), attributed to data scarcity.
-Physics coupling studies indicate compressible flows demand tripled resources due to interdependencies, and LLM bias toward incompressible setups causes dimensional errors in complex cases.
-These findings highlight ChatCFD’s robustness, efficiency, and cost-effectiveness, establishing benchmarks for AI-driven CFD.
+By integrating a multi-agent architecture, domain-specific structured knowledge bases, precise error locator, and iterative reflection, ChatCFD dramatically outperforms prior systems while exposing fundamental limitations of current LLMs in scientific computing.
+On 315 OpenFOAM benchmark cases, ChatCFD attains an 82.1% execution success rate—far surpassing MetaOpenFOAM (6.2%) and Foam-Agent (42.3%).
+More importantly, we introduced physical fidelity as a rigorous new metric: among all attempts, ChatCFD generates physically correct solutions in 68.12% of cases, with 60% of semantic failures traced to subtle boundary/initial condition errors that escape convergence checks.
+A dedicated Physics Interpreter further achieves 97.4% summary fidelity, revealing the striking gap between LLMs’ fluency in high-level narration and their struggle to enforce dozens of tightly coupled domain constraints in executable code.
+Resource-efficiency analysis on 205 benchmark cases further underscores ChatCFD’s practical superiority.
+It consumes only 192.1k tokens and \$0.208 per case—roughly half the tokens of Foam-Agent and 1.5× cheaper than MetaOpenFOAM—while requiring the fewest reflection iterations on average.
+This efficiency gain over naive LLM prompting stems from strategic model routing (DeepSeek-V3 for generation, DeepSeek-R1 for reasoning) and a reflection memory that filters irrelevant context.
+Even under a strict iteration limit, ChatCFD’s reflection trajectory consistently outperforms baselines, solving cases that saturate the others at  50% success after 25 iterations.
+These results demonstrate that structured domain integration not only boosts accuracy but dramatically reduces the operational cost of LLM-driven scientific computing, making high-fidelity CFD automation viable at scale.
+Ablation studies confirm that structured OpenFOAM knowledge is indispensable: removing the Solver Template DB alone collapses accuracy to 48%, while the Error Locator Module proves the single most critical component for real-world robustness.
+Flexibility experiments demonstrate ChatCFD’s ability to autonomously select appropriate solvers across compressible/incompressible and steady/transient regimes (95.23% success) and switch turbulence closures (100% success), even on unseen configurations.
+By reproducing complex literature cases—NACA0012 airfoil, supersonic nozzles—ChatCFD shows 60–80% end-to-end success where baselines fail entirely.
+These results establish potentially new benchmarks for AI-driven CFD and provide the quantitatively grounded evidence of where current LLMs break in scientific tool use.
 ChatCFD effectively addresses the critical gaps in CFD automation agents.
@@
 Third, by leveraging extensive literature-derived corpora with detailed specifications and benchmarks, it will facilitate scalable testing and iterative agent evolution, harnessing millions of CFD papers for refinement and validation in the future.
-These advancements accelerate scientific discovery in fluid mechanics and engineering by automating hypothesis testing, parameter optimization, and multiphysics simulations.
+These advances will enable scientists across disciplines to explore bold ideas using CFD that were previously stalled by technical details.
+By liberating researchers from low-level case-setup burdens, it enables non-experts to perform computational fluid dynamics studies, allows rapid exploration of hypothesis and design spaces, and frees CFD specialists to focus on genuine scientific questions.
 ACKNOWLEDGEMENTS
@@
 Author Contributions
-E Fan: Methodology, Software, Data curation, Formal analysis, Resources, Writing - original draft.
-Kang Hu: Methodology, Software, Data curation, Formal analysis, Resources.
+E Fan: Methodology, Data curation, Formal analysis, Writing - original draft.
+Kang Hu: Methodology, Software, Data curation, Formal analysis, Writing - original draft.
 Zhuowen Wu: Methodology.
@@
 [11]
-Figures 12 and 13 show the velocity magnitude contours and surface pressure coefficient distributions for the NACA0012 airfoil Cases 1 to 4 at 10° angle of attack using different turbulence models.
+Figures 17 and 18 show the velocity magnitude contours and surface pressure coefficient distributions for the NACA0012 airfoil Cases 1 to 4 at 10° angle of attack using different turbulence models.
 As can be seen, the velocity contours from the four cases are similar, and the pressure coefficient distributions agree well with the experimental results [11].